### v1

# Setting the Course, but Forgetting to Steer: Analyzing Compliance with GDPR’s Right of Access to Data by Instagram, TikTok, and YouTube

Abstract
The comprehensibility and reliability of data download packages (DDPs) provided under the General Data Protection Regulation’s (GDPR) right of access are vital for both individuals and researchers. These DDPs enable users to understand and control their personal data, yet issues like complexity and incomplete information often limit their utility. Also, despite their growing use in research to study emerging online phenomena, little attention has been given to systematically assessing the reliability and comprehensibility of DDPs.
To bridge this research gap, in this work, we perform a comparative analysis to assess the comprehensibility and reliability of DDPs provided by three major social media platforms, namely, TikTok, Instagram, and YouTube. By recruiting 400 participants across four countries, we assess the comprehensibility of DDPs across various requirements, including conciseness, transparency, intelligibility, and clear and plain language. Also, by leveraging automated bots and user-donated DDPs, we evaluate the reliability of DDPs across the three platforms. Among other things, we find notable differences across the three platforms in the data categories included in DDPs, inconsistencies in adherence to the GDPR requirements, and gaps in the reliability of the DDPs across platforms. Finally, using large language models, we demonstrate the feasibility of easily providing more comprehensible DDPs.
  *These authors contributed equally to this work.
1 Introduction
Ensuring compliance with the General Data Protection Regulation (GDPR) [12] is essential for platforms handling personal data, as it establishes legal requirements for data processing, security, and user rights. One of these rights is the right of access by data subjects, which enlists provisions whereby data subjects (i.e., end users) can request access to their data being processed by platforms (Article 15(3) GDPR). Platforms have to comply with these requests within a stipulated period of one month and provide the user with a copy of their data. Usually, platforms provide this data in the form of data download packages (DDPs) [34]. These DDPs often include information about the users’ activities on and off the platform, device and app information, inferred preferences, and other information depending on the business of the platform.
Implementation across platforms: In fact, the GDPR stipulates that information shall be provided to users in commonly used electronic forms (Art. 15(3)(2) GDPR). Such an implementation is crucial for enabling users to exercise their right to data portability (Article 20), which allows end users to transmit data from one platform to the other without hindrance [12]. However, the GDPR does not provide explicit standards for the structure or content of DDPs, leading to varying implementations of Article 15 across platforms [8].
Understandably, platforms that offer different types of services interpret and implement the GDPR provisions differently. For instance, the DDP of a digital marketplace like Amazon is naturally distinct from that of a digital public space like TikTok. However, platforms offering similar services should be able to provide comparable information to users. For example, platforms providing short-format video streaming services should, in principle, offer similar details through their DDPs. To the best of our knowledge, such a comparative study of the implementation of the right of access under the GDPR is rare in the literature. This observation leads to a key unexplored research question as follows: RQ1: How similarly do platforms offering similar services implement the right of access under the GDPR (Article 15(3))?
Desideratum 1: Comprehensibility: Irrespective of the answer to RQ1, these implementations (DDPs) should adhere to a set of requirements outlined in the GDPR to ensure user comprehensibility. According to Article 12(1), these DDPs should be (a) concise, (b) transparent, (c) intelligible, (d) easily accessible, and (e) using clear and plain language [12]. However, neither the GDPR nor accompanying guidelines [14] provide a clear technical definition of these criteria. Further, adherence to these requirements in the current implementations also remains unexplored in the research community. The authors in [8] came closest to understanding the question of adherence, but they restricted their analyses to conciseness. This gap leads us to the second research question in this paper: RQ2: How comprehensible are the DDPs provided by platforms under Articles 12 and 15 of the GDPR?
Desideratum 2: Reliability: Finally, another important expectation for the effective implementation of the provision is that the information in the DDPs is reliable. With the increased restriction on access to platform data for research [33, 4], many researchers are relying on data elicited through exercising the right of access for understanding the dynamic interplay between population and platforms. In these so-called “data donations” based studies, consenting participants first request their own data from the platforms, and then provide this data to researchers, typically for financial compensation. Several recent scientific studies use this setup to explore fundamental questions about digital platforms— such as user engagement [38], content recommendation [35], and ad targeting processes [36] — and even questions beyond online activities, e.g., offline risks [3] or the likelihood of addiction [37] using the resultant DDPs. However, if the reliability of the information provided by the platforms is compromised, then the insights drawn from such data sources might be invalid.
To this end, a recently conducted survey among researchers indicates that 86% of researchers report concerns regarding the reliability of the data collected using DDPs [34]. The most important dimensions of reliability that researchers are concerned about include: (a) completeness, (b) correctness, and (c) consistency of the shared information. This brings us to our final research question: RQ3: How reliable is the information within the DDPs provided by platforms under Article 15(3) of the GDPR?
The current work: Although the current work could have been conducted on any digital platform, due to the growing popularity in consumption of short–format videos [38, 25] and the impact they have on the population [13], we investigate the aforementioned research questions on Instagram, TikTok, and YouTube.  Figure 1 describes the pipeline of our conducted research. After analyzing the DDPs shared by these platforms, we conducted an extensive survey among 400 participants from Germany, France, Italy, and Spain to evaluate their assessment of the comprehensibility of the respective DDPs. Further, by utilizing (a) participants’ and (b) European Data Protection Board’s interpretations of the requirements, (c) current advancements in large language models (LLMs), and (d) our technical expertise, we propose data representations for the different categories of information. These recommended representations are then evaluated against the best representations of respective categories in the current implementation of the platforms by 200 participants.
To answer the questions regarding the correctness and completeness of the DDPs, we browse the three platforms using sock-puppet accounts to (a) automatically log the browsing behavior and (b) request the DDPs from each of the platforms. We repeat this process for a period of one month to understand the consistency in the shared DDPs for the accounts. In addition, to better understand the consistency across various accounts, we also collect DDPs requested by real-world participants from the three platforms.
Figure 1: Our pipeline to evaluate comprehensibility and reliability of the current evaluation of Article 15(3) of the GDPR.
We summarize our major findings with respect to each research question below.
RQ1: The studied platforms implement Article 15 of the GDPR differently: Instagram, TikTok, and Youtube share different categories of data in their DDPs. While Instagram and TikTok share more data with their users, YouTube’s shared data is rather limited. Irrespective of the amount of data, the only underlying similarity is that none of the platforms share the purpose of data collection and processing in their DDPs.
RQ2: Instagram’s current data representation is the most comprehensible: Based on the preference votes of 400 participants, we find Instagram’s DDP adheres to GDPR’s requirements more than the other two platforms.
RQ2: Considering interpretation from different stakeholders significantly improves the data representation: 200 survey participants prefer our proposed data representation to current representations along all requirements except conciseness. In fact, our proposal is preferred in as many as 44 out of the 52 evaluations (13 categories $\times$ 4 requirements).
RQ3: TikTok’s DDP is the most reliable as per our evaluation: Our analyses using sock-puppet accounts reveal TikTok’s DDP presents a complete, correct, and consistent representation of our browsing behavior. Instagram and YouTube exhibit varying degrees of missing data across DDPs collected over the period of our study. However, the analyses among real users show Instagram and TikTok share data for different periods in their DDPs for different categories, whereas YouTube shares all the data types for similar duration.
To the best of our knowledge, this study is the first detailed evaluation of adherence of different requirements mentioned in the GDPR. The insights drawn from our study further underline the need for standardization and follow-up on the implementations to ensure that the primary goals of data protection laws are achieved.
2 Background and Related Work
To position the current work in the perspective of the literature, we present a brief overview of the following: (a) GDPR and the rights of access, (b) leveraging data download packages for platform audits, (c) audits of implementations of provisions of the GDPR.
GDPR and the right of access: The GDPR is a landmark privacy law designed to give individuals greater control over their personal data and to standardize data protection rules across the European Union (EU) [12]. While the GDPR has many provisions and rights enlisted, our work focuses on Article 15(3): the right of access, which allows users to request a copy of their data collected by online platforms. Most platforms implement this requirement by providing a Data Download Package (DDP) upon request. For such implementations, GDPR provides a set of desiderata for the DDPs. Specifically, Article 12(1) requires a DDP to be concise, intelligible, transparent, accessible, and in clear and plain language [12]. Moreover, the GDPR gives the right to end-users to be able to port their data from one platform to another under Article 20. However, if platforms offering similar services implement these rights in significantly different ways, platforms will essentially deprive end users of their right to portability.
Leveraging data download packages: The richness of DDPs has enabled new research directions previously deemed infeasible. Indeed, many recent studies are utilizing them through donations made from participants with their explicit consent. To this extent, researchers have used DDPs to conduct research on topics related to personal health and safety [20, 37, 3], news and politics [6, 17, 18], auditing recommendation algorithms [24, 35], analyzing user behavior on social media platforms [38, 15], and understanding ad targeting types and user perceptions [36].
Audits of implementations of GDPR: Many studies have investigated (a) whether consent forms are designed properly by platforms [29, 30, 23, 9], (b) whether platforms are undermining user privacy under the pretense of ‘legitimate interest’ [31, 21, 22], (c) inconsistencies in platform’s data withdrawal behavior [11]. However, little attention has been given to platforms’ implementation of the right to access. A recent study shows that end users often view DDPs as overwhelming and unclear, leaving key questions unanswered while triggering privacy concerns [8]. Moreover, researchers have expressed concerns about the reliability of the DDPs and have faced challenges in using them [34, 19].
The current work: To the best of our knowledge, our work is one of the first studies that tries to unravel the current implementations of Article 15(3) of the GDPR from the lens of comprehensibility and reliability. Furthermore, our proposed recommendations for the potential implementation of DDPs may not only improve the current implementations, but also help in the implementation of other rights provided to users under the GDPR, such as the right of data portability.
3 RQ1: How similarly do platforms offering similar services implement GDPR Article 15?
Before analyzing the comprehensibility and reliability of DDPs, we first look into the content of the DDPs provided by each of the platforms (i.e., we compare how the platforms are currently implementing GDPR Article 15(3)).
3.1 Platforms under consideration
Since the primary focus of the study is to understand how similar platforms implement the same provision in the GDPR, we decided to focus on multiple platforms providing similar services. To this end, there are many options for digital streaming platforms (e.g., Netflix, Disney, etc.), digital marketplaces (e.g., Amazon, Zalando, etc.), and social networking platforms (e.g., Facebook, X, LinkedIn, etc.). In this study, we focus on three platforms that provide short-format video streaming services: TikTok, Instagram, and YouTube. Our motivation to study these platforms stems from the recent increase in popularity of short-format video platforms [38, 25]. Furthermore, all three platforms have been designated as Very Large Online Platforms (VLOPs) by the EU under the recently enacted Digital Services Act (DSA), indicating the far-reaching impact that these platforms have on the European society [13].
TikTok, Instagram, and YouTube implement Article 15(3) of the GDPR, through which they share DDPs with their end-users upon request. In the current implementations, platforms share data primarily in two formats (a) machine-readable – which is usually a single JSON or a collection of JSON files; (b) human-readable – where Instagram and YouTube opt for HTML version, whereas TikTok shares it in TXT files segregated across many directories. If one takes a closer look at the DDPs received from any of the platforms, their contents can be broadly divided into the following categories: (a) user’s usage, (b) user’s content, (c) personal details, (d) advertisements, (e) miscellaneous. Next, we examine the DDPs that Instagram, TikTok, and YouTube provide and observe their content and presentation. Most of the observations discussed below are summarized in Table 5 (see Appendix A).
3.2 Information regarding users’ usage
One of the most important facets of the information contained within the DDPs is how a user uses the social media platform. Such usage information may include the content that a user watches, searches for, or engages with implicitly or explicitly.
Watch history: Watch history is an ordered list of all the content that a user has watched before requesting the data. DDPs of all three platforms contain watch history data of the data subject. However, as Table 1 shows the data provided across the three platforms are widely varying. Among these, two points are strikingly different – (a) while YouTube provides this data for the entire lifetime of a user on the platform, Instagram provides data of at most the last two weeks and TikTok provides data for around six months before the data request (without any clear explanation for the reduced duration); (b) while TikTok and YouTube provide the watch history all at once, Instagram segregates it into three files based on the type of content, i.e., ads, posts, and videos watched by the user.
Feature YouTube Instagram TikTok
Duration Entire user’s lifetime 2 weeks 6 months
Content URL
Timestamp
Content title
Author name (with author’s URL)
Ad identification (whether ad or not)
File segmentation Single file
Three files
Single file
Table 1: Comparison of watch history data across platforms.
Like history: Similar to watch history, like history is an ordered list of all the contents that a user has liked before requesting the data. The details across the three platforms are noted in Table 2. The table demonstrates that there is a clear lack of consistency in how platforms currently implement Article 15(3) GDPR. While one platform (YouTube) records the like history on its application, it does not share this information in the YouTube DDP  [^1] Although Like history is not there in YouTube ddp, it is included in the Google DDP.. Another platform (Instagram) provides the watch history for just a couple of weeks without any link to the watched content, but it provides the like history of a lifetime, including references to the liked content.
Platform
Provides
like
history?
Details provided
TikTok
Link to the liked content,
timestamp of when it was liked
Instagram
Link to the liked content, author details,
timestamp of when it was liked
YouTube
Shows like history on mobile app and web
version but does not share it in the DDP
Table 2: Comparison of like history details across platforms.
Time spent on platform: While all three platforms have a detailed dashboard that mentions the average time spent and the exact amount of time users spend on the application daily, the same information is not provided in the DDP in response to Article 15 GDPR. On a positive note, TikTok recently added a new field in the data, namely “Activity Summary,” which enlists how many videos a user has commented on, shared, or watched until the end since their registration.
Other usage activities on platform:  Table 5 (in Appendix A) shows some other important usage activities. Most of the activities, e.g., comment, search, save, share, and writing a message are recorded by all platforms where the features are applicable. However, sharing across applications, i.e., when a user shares the video on other social media platforms or copies the link for posting elsewhere, is only shared in TikTok DDPs. Also, while both TikTok and Instagram maintain a list of inferred interests of the data subject, YouTube DDPs do not have them.
3.3 Information regarding user’s content
While the usage data is about how users consume or behave on video streaming platforms, information regarding their content refers to the content that a user creates and uploads to the platform for others to consume. This information encompasses the media (image/audio/video) a user uploads on their profile, along with any textual captions, locations, date and time information, etc.
Aspect Details provided Platforms
Content created
Media (image/audio/video),
textual captions,
date, and time information
TikTok,
Instagram,
YouTube
Addnl. details
Metadata such as software used
for uploading (e.g., Android gallery),
device ID, camera metadata
Instagram
Location details
Longitude and latitude of upload site
if the author tags the location
TikTok,
Instagram,
YouTube
Table 3: Details of user created content shared by platforms.
While all three platforms provide a copy of the media and text details, Instagram shares additional information including software used to upload the content, device ID, metadata about the camera, etc. (Table 3).
3.4 Information regarding personal details
While the two facets discussed above focus more on the content that users consume, engage with, or create, the information regarding personal details is more privacy-sensitive. These details contain basic personally identifiable information (PII), e.g., account details (which may include one’s name, phone number, e-mail address, date of birth, profile picture, etc.). However, apart from these basic personal details, most of the DDPs also include information about the different devices through which people visit the platforms. TikTok and Instagram directly store these details as part of the login information, whereas YouTube does not offer this data in its DDP. For YouTube, this data can be accessed only if someone requests their entire Google DDP. [^2] One justification may be that one can not sign in to their YouTube account unless they sign in to their Google account on the same device.
Login history: In light of the prior descriptions, login history contains the list of login activities made by the data subject. The login history details are significantly different in TikTok and Instagram DDPs. TikTok provides the timestamp, IP, device model, operating system, network type, and carrier provider in its DDP. On the other hand, Instagram’s login history data is very detailed. Along with providing the above details, Instagram provides (and therefore collects) cookie information, language code, Instagram app version, display properties of the device, hardware identifier, and some internal identifiers. Such differences in data collection and sharing practices point to two important possibilities: (a) TikTok might be collecting some of these details and not sharing them with its end-users and, therefore, is not being transparent; (b) Instagram might be collecting more information than necessary, thus becoming more intrusive to privacy. Diving deep into these possibilities is beyond the scope of this study, and we would like to explore these directions in future work.
3.5 Information regarding advertisements
Advertisement on social media platforms is a prevalent phenomenon, as users are constantly served with ads when they scroll through Instagram, TikTok, or YouTube. However, based on our observations, there is no standardized way of including ads that users have seen in their requested DDP.
Ads viewed: As mentioned in Section 3.2, Instagram maintains a separate list of ads viewed by users. However, the data provided by Instagram only mentions the author/advertiser name and timestamp. TikTok, on the other hand, does not demarcate any advertised content that a user has seen in their timeline. In contrast, YouTube distinguishes all the ads that a user has encountered from the organic content. Also, YouTube provides a link to the ad video along with its title and when it was viewed. However, neither of the three platforms provides information on ad targeting parameters. From the ads viewed details, a user can understand these are the ads they got to watch on their timeline, but why those particular ads were served to them remains an enigma. While this information does appear when an advertisement or paid content appears on the three platforms, it is not included in the shared DDPs.
Ad targeting data sources: The Instagram DDP contains a list of advertisers who have used data about one’s online activity or profile for targeted advertising purposes. This list also conveys information about how these advertisers obtained users’ data. In specific, as per this information, there are three ways in which the advertiser might target a person: (a) custom audience – where the advertiser targets the data subject through a custom audience with a list of customer data (e.g., email addresses, phone numbers, or other identifiers) which the advertiser might have obtained from other third party sources; (b) remarketing custom audience – where the advertiser targets the data subject because (s)he might have interacted with some of the advertiser’s content, visited their website, etc.; (c)  in-person store visit – where advertiser can use location data or check in data because the data subject might have visited one of their physical stores. The DDPs of YouTube and TikTok do not provide any such information on advertisers and how they gather data about a particular user.
Off-platform activities: Another important source of data through which these platforms collect enormous amounts of digital behavioral data is off-platform activities. This is a list of all the activities in external platforms (outside the source platform) that these platforms have tracked or linked to a user. Note that this might be an artifact of the said external platforms’ installed Meta/Google/TikTok pixels. A user might have used a Meta or Google login to sign up, other tools embedded in the software development kit of the external platform, or any data-sharing agreement between the external company and any of the platforms we investigate.
For instance, if a user visits the website or application that also installs one of the tracking techniques of any of the short-format platforms, then the said website can share the user’s activities, such as page view, purchase, search, logins, etc., along with other activities, such as adding a product to the cart, view contents, install app, launch app, etc. Instagram and TikTok DDPs contain off-platform activities of users, whereas YouTube does not have such data. Both Instagram and TikTok share the external platform name, event or activity type, and the timestamp in their DDPs.
The DDPs from the three platforms also provide a wide variety of other details. For example, they may contain information on link history, survey, application setting, shopping, etc. For more details on this, please refer to the summary of all the discussions above, listed in Table 5 (see Appendix A). All three platforms miss one of the important goals set by GDPR Article 15(1)(a), which is to include the purposes of the data processing and collection. None of the DDPs include a purpose statement explaining why certain data is collected, nor do they provide a README or manual to help users understand the significance of the data being collected.
Important takeaways
[…]
  The three platforms implement Article 15 very differently. The Instagram and TikTok DDPs contain a significantly larger number of data types than the YouTube DDP. Instagram shares (and thus collects) more data.
[…]
4 RQ2(a): How comprehensible are DDPs?
Here we investigate whether the current implementations of platforms adhere to the requirements Chapter 3 of the GDPR outlines a number of rights of data subjects (i.e., end users), including the right of access (Article 15) [12]. Before outlining these rights, the GDPR provides a list of requirements for how information should be communicated to end users. Specifically, the first paragraph of Article 12 of GDPR reads, “The controller shall take appropriate measures to provide any information referred to in Articles 13 and 14 and any communication under Article 15 to 22 and 34 relating to processing to the data subject in a concise, transparent, intelligible and easily accessible form, using clear and plain language $\dots$ [12].
As the GDPR does not provide more succinct definitions of these requirements, we adopt two approaches to interpret these terms. First, following trends in empirical contract research to employ user surveys for interpretation of legal terms [5], we conduct a large-scale user survey among participants from different European countries to understand their interpretations of some of the requirements. Second, we apply the interpretations of the requirements by the European Data Protection Board (EDPB) [14]. Finally, we evaluate the current implementations, i.e., the DDPs obtained from Instagram, TikTok, and YouTube, against the interpretations by users and the EDPB.
Participant recruitment: We recruited participants to conduct our survey from the crowd-sourcing platform Prolific[27]. On Prolific, we hired 100 participants each from Germany, France, Spain, and Italy (400 in total) who had a high approval rate of at least 98% to take part in our survey. Our motivation for a general set of respondents stems from the fact that the GDPR entitles general residents in the European Union(EU) with the right to data access, and requirements are intended to benefit these residents. Though, eventually, our plan is to extend the study to all countries within the EU, we chose to start with the countries that have the highest GDP [32]. We also selected the standard gender breakdown while setting up the surveys. In total, we recruited 228 males (57.0%), 161 females (40.3%), 9 self-reported as “other” (2.3%), and 2 participants who preferred not to disclose their gender (0.5%) (See Table 6 in Appendix C for details).
Prior to entering our survey, we presented an online consent form for expressing explicit consent from the participants. Upon completion of the survey, we compensated each participant at a rate of £9 per hour, which is recommended by Prolific to be a good and ethical rate of remuneration [26]. Our survey comprised three components: (a) Awareness of participants about the rights, (b) Interpretation of the requirements by the participants and (c) adherence evaluation. Next, we elaborate abut the survey setup and important observations for each of the components.
4.1 Awareness about the GDPR right of access
Survey setup: In the first component, we first tried to educate our participants about Article 15 of the GDPR and how they can request their data from online platforms. We also provided a link to a Google Drive folder where they could see an example of a DDP. At this stage, we asked our participants three questions regarding (1) whether they were aware of such right before taking our survey, (2) whether they had exercised this right by requesting their data, and (3) if they had exercised their rights, then why. We designed this part of the survey to understand the general awareness of the participants and make them aware of the said right.
Observations: Out of the 400 participants, nearly $72\%$ participants responded that they were aware of the right of access to data before participating in our survey. At the same time, only $29.2\%$ of them answered in affirmation when it comes to exercising their rights by requesting the data on some platform. Although there is a massive gap between awareness and exercise of the data requests, the numbers are surprisingly high. We also asked our participants to provide reasons for exercising their data requests in free-form tex. Through manual annotations the responses were characterized into four groups. Almost half of the participants (51%) mentioned curiosity, knowing what information platforms collect about them, to be the primary reason for their data request. Further, 18% mentioned seeking specific information, e.g., identifying a song they had listened to earlier, cross-checking specific details, or determining the amount of time spent on the platform, etc., as the primary reason for their requests. Finally, 11% of our participants made their requests to keep a backup.
Interestingly, one-fifth of the participants (20%) mentioned that they requested data from several platforms for participating in some research study. Note that several recent studies–for understanding social media– use data donations from users on Prolific ([37], [35]) and other crowd-sourcing platforms as their primary data source. This observation likely explains the surprisingly high percentage of awareness and exercise statistics that we reported above.
4.2 Interpretation of the requirements
Survey setup: To understand the interpretation of requirements by common end users, in the second part of our survey, we showed the participants the paragraph of Article 12 (see above), which mentions the different requirements of DDPs. We asked our participants about their interpretation or expectation of the shared personal data in the DDPs to follow any of these requirements. Next, we elaborate on the different interpretations of these requirements from the responses of our participants while contextualizing them further with the guidelines of the EDPB [14], which represents the data protection authorities of all EU member states.
Conciseness Interpretations by the survey participants: We perform thematic analysis of the responses provided by the 400 participants using BERTopic [16]. The important themes that resonated across the responses spanning all countries are – prioritize important & relevant data, summarize important observations, and present data in a structured way that is easier to navigate.
However, a fourth theme emerged in the interpretations shared by participants from Italy, France, and Spain, but not by those from Germany. Specifically, these participants touched upon the concept of data minimization. To this end, some participants mentioned the documents should be shorter, while others mentioned avoiding an overwhelming amount of text. In short, while participants from Germany only prioritize the informativeness of the content, other participants prioritize the length of the content as well.
Interpretations by the EDPB: In this context, the European Data Protection Board interprets concise presentation to be done efficiently and succinctly to avoid information fatigue [14]. While the above interpretation falls within the interpretations mentioned by our survey participants, the EDPB also asserts that the presentation should differentiate privacy-sensitive information from non-privacy-sensitive information.
Transparency Interpretation by the survey participants: Unlike the case of conciseness, there are no major differences in how participants from all countries interpret transparency. The interpretation of transparency can be divided into two major aspects. First, participants call for unambiguous, full disclosure of all the data that has been collected about them in the way they are collected. Second, and most importantly, some participants also expect the DDPs to mention how the data is processed and for what purpose the data is collected. To this extent, the response from one of the participants captures all the notions beautifully: “It (Transparency) means providing clear honest and complete information about how the data is being used. There should be no hidden details or surprises for the individual. People should easily understand what data is being collected why its being collected who will access it and how long it will be kept.” At this point, we would also like to note that as per the participants’ interpretation, conciseness and transparency are two competing requirements, i.e., to satisfy one, the other possibly has to be violated.
Interpretation by the EDPB: While the EDPB draws most of its interpretations from Recital 39 of the GDPR, it interprets transparency as allowing data subjects to determine the scope and consequences of the concerned data processing [14]. To this end, the EDPB calls for data controllers (i.e., platforms) to separately spell out the most important consequences and risks that the data processing may entail.
Intelligibility Interpretation by the survey participants: Among all the requirements provided by the GDPR, the most subjective one is that of intelligibility. However, almost all the participants explicitly interpret it to be easy to understand and link it closely with the language being straightforward and without any technical and/or legal jargon. Participants from all countries also mentioned that intelligible data should have a clear explanation of how and why a certain piece of information is collected. The participants also noted some interesting expectations in this regard. For instance, participants from Germany mentioned that platforms shall share their DDPs in one’s language of comfort (native language or other language of choice). Moreover, participants from Italy stated that platforms should share their DDPs in the form of elegant visualization.
Interpretation by the EDPB: The EDPB describes intelligibility to be a representation that should be understood by an average member of the intended audience. In addition, they also mention that it is closely linked to the requirement to use clear and plain language.
Clear and plain language Interpretation by the survey participants: Understandably, most of the participants interpret plain and simple language to be words that are easily understandable and unambiguous. Some participants elaborate on the nuance by stating that legal and technical terms should be avoided or at least explained in a simpler way. To this end, one clear observation from our survey responses is that participants do not distinguish between intelligibility and language.
Interpretation by the EDPB: Along the lines of the interpretations mentioned above, the EDPB interprets this requirement as being definitive and should not be phrased in ambivalent terms or leave room for different interpretations [14].
Accessibility Interpretation by the survey participants: Most of our participants interpret accessibility as being easily available, i.e., it is easy to find the place to request the data and to access the data. However, there are certain interesting and valid interpretations which are worth mentioning. Some participants mentioned that the data needs to be available in a format so that it can be opened without any sophisticated software or tools. At the same time, the timeliness at which the data is provided upon request is another interesting aspect touched upon by some of our participants. Another interesting interpretation was the data should be available to everyone, including people with visual impairments by using accessible design features like screen readers or alternative text making sure the users can access data without any barriers.
Interpretation by the EDPB: The EDPB’s interpretation of accessibility also follows from the above popular responses from the survey participants, wherein they say that it should be immediately apparent to end-users as to where and how the DDP can be accessed [14].
Our observations on accessibility: Despite agreeing with all of the above interpretations, we observe a few differences across the platforms because accessibility is the most objective property for evaluation. In terms of ease of request, from the content page of a user’s account, one needs to click 6, 6, and 10 times on TikTok, YouTube, and Instagram, respectively, to be able to request the data in the human-readable format. Notice that, while on Instagram, a user needs to click more, the platform provides many options to the end-user, such as the time duration that their DDP should contain. Moreover, based on our anecdotal observations, the turnaround time for TikTok upon request was instantaneous during this research study. For Instagram and YouTube, it often takes 10 to 15 minutes for the DDP to become available. However, all the turnaround times are within the GDPR-prescribed maximum duration of one month. Further, one can download one’s data on TikTok and Instagram from both their applications and web versions, but YouTube shares the DDP with its end users by sending an e-mail with a URL to the DDP. All these platforms provide DDPs in a compressed format, which makes it harder for the end-users to open them on a smartphone. At the same time, an end user needs to have a browser or text reader to open the HTML (YouTube and Instagram) and TXT (TikTok) DDPs. In summary, while the data request and download are relatively accessible in the current setting, reading the data is still complicated for the average end user. At this point, we would also like to note that we are not aware of any special provisions that platforms may have for people with visual impairments to access the data.
4.3 Adherence evaluation
Survey setup: In the final part of the survey, we presented our participants with the information shared in the DDPs of Instagram, TikTok, and YouTube in an anonymized and randomized format. We showed different categories of content and the whole DDP to our participants. The different categories of content that we shared include data from all of the categories of information that we elaborated in Section 3.  Table 7 (refer Section C.1) summarizes the specific thirteen categories that we evaluated. [^3] We should point out that YouTube does not provide like history, autofill, location, and off-platform activity. Hence, for these categories, we asked our participants to evaluate between Instagram and TikTok. Note that, for this comparative study, we used DDPs that are shared in a human-readable format by all three platforms. Instagram and Tiktok share the human-readable version in HTML and TXT formats, respectively. However, YouTube DDP only contains watch history, search history, personal information, and login history in HTML format. All other categories of data are found in CSV or JSON formats. Since the primary goal of this study is to compare the contents, we showed them consistently in a $<category,value>$ pair in our survey. We acknowledge that the way data is shown to users may affect some of the results. However, in this study, we restrict ourselves to the evaluation of what content is shared rather than how it is shown. We plan to explore the visualization aspects in future studies.
For each category, we showed participants the content shared by the three platforms side by side and asked them which platform’s data representation they found to be more concise/intelligible/transparent/in clear and plain language. [^4] We exclude accessibility from the evaluation because it pertains to getting access to request and download DDPs more than to their representations. Also, we stated our observations about this requirement in Section 4.2.  Figure 8 (see Section C.1) depicts a representative example of one of the questions that we used in our survey. The first preference in Figure 8 means the most preferred choice of data representation of a participant, whereas the third preference means their least preferred choice of data representation. We posed this question as a multiple-choice grid, so each participant had to provide an ordered preference for the three representations. For example, if a participant rates Instagram’s representation as their first preference, then they cannot rate any of the other platforms as their first preference.
Note that the final part of the survey consisted of 14 (13 categories + overall DDP) $\times$4 evaluations where participants gave us an ordered preference. Hence, to avoid participant fatigue, we asked 50% of our participants to evaluate intelligibility and transparency and asked the other half to evaluate conciseness and language properties. Hence, 200 participants answered each of the evaluation questions.
Observations
Next, we report our findings on the adherence evaluation of four requirements.
Language: Figure 2 shows the percentage of first preference votes elicited by the representation of each of the platforms for the four requirements that we surveyed. In the evaluation of clarity and simplicity of language, Instagram won the first preference votes in eleven out of the thirteen data categories (Figure 2(a)). In contrast, for the remaining two categories, participants preferred the language in which YouTube presents its search and watch histories.
To understand the overall preference of participants across the board, we weight every first preference vote with a score of 3, second preference vote with a score of 2, and third preference vote with a score of 1. For categories, where we are comparing only two platforms, we assigned scores 3, 1 for first, second preference respectively. Then, we evaluate an average preference score for each of the platforms based on participants’ preference orderings. Figure 3 shows the average preference scores of each of the platforms in each category of data. Based on the average preference scores, apart from watch, search histories and location, the gap between preference toward the language of Instagram’s data representation is way more than that of TikTok and YouTube (Figure 3(a)). Across the different categories of data, the participants’ average preference score for Instagram’s language is 2.48, whereas TikTok and YouTube fail to reach an average preference score of 2.
Apart from the thirteen categories of data, we also asked our participants to evaluate the overall DDP from the three platforms. When participants were asked to evaluate the overall DDP, 55% of them preferred the language of Instagram’s DDP over the other two. TikTok and YouTube DDPs are preferred by 25% and 20% of the participants, respectively.
Intelligibility: Similar to Language, Instagram won the first preference vote for eleven out of the thirteen categories evaluated (Figure 2(b)). Among these, except for devices and location categories, nearly more than 75% of the participants preferred Instagram’s data representation to be the most intelligible.
The average preference scores almost follow the same trend as that of language, where the Instagram DDP’s intelligibility (2.53) is substantially higher than that of YouTube (1.75) and TikTok’s (1.64) (Figure 3(b)). For the evaluation of the overall DDP, 63% of the participants identified the Instagram DDP to be the most intelligible. TikTok and YouTube were chosen as the first preference by a considerably smaller proportion of our participants, nearly 18% each.
Transparency: YouTube won the first preference for five categories, where over 70% of the participants chose it for watch history, search history, and login, followed by devices and content with 55% and 46%, respectively (Figure 2(c)). For the remaining eight out of the thirteen categories, Instagram won the first preference of the participants in a large percentages.
We find that the average preference scores of Instagram and YouTube are comparable (Instagram: 2.23, YouTube: 2.15), whereas TikTok (1.66) has lower score (Figure 3(c)). In the evaluation of the overall DDP, Instagram is preferred more compared to the other two platforms, with nearly 46% of the participants choosing it to be the most transparent. On the other hand, YouTube is preferred by 35% of the participants, and the remaining 19% selected TikTok.
Conciseness: Based on Figure 2(d), we found that in nine out of thirteen categories of data that we surveyed, respondents preferred TikTok’s data representation as the most concise, whereas Instagram’s data representation is found to be the most preferred for the remaining four categories. While TikTok’s representation is preferred substantially in the location, autofill, and off-platform categories ($\geq$60first preference votes), Instagram’s representation is preferred more for likes history and contents ($\geq$65first preference votes). Based on the average preference scores for conciseness, we see that the representations are found to be very comparable in most of the data categories (Figure 3(d)). The average preference scores of Instagram and TikTok are very comparable (Instagram: 2.12, TikTok: 2.15), whereas that of YouTube is considerably lower (1.60). While evaluating the overall DDP, 43% of the participants found Instagram’s DDP to be more concise, followed by Tiktok and YouTube with 35% and 22%, respectively.
Variations across countries: We do not observe qualitative variation across the four countries for three requirements – language, intelligibility, and transparency. However, for conciseness, participants from Germany predominantly preferred Instagram, with over 50% of the participants selecting it as their first choice across most categories. In contrast, participants from other countries primarily selected TikTok as having the most concise representation for all the categories, except likes, connections, and content. To understand this, we manually examined how participants across different demographics interpret conciseness. We found that, except for participants from Germany, most demographics have an interpretation of conciseness as: (1) minimizing document length, (2) avoiding information overload etc. This difference in interpretation is reflected in how our participants evaluated the data representations: while participants from Germany gave more priority to content, other participants gave more priority to data categories having less content.
(a) Clear and Plain Language
(b) Intelligible
(c) Transparency
(d) Conciseness
Figure 2: Percentage of first preferences across all categories and for the entire DDP. Instagram is the top choice for most categories across all requirements, except for conciseness. For watch and search histories, YouTube is the most preferred platform in transparency, intelligibility, and language.
(a) Clear and Plain Language
(b) Intelligible
(c) Transparency
(d) Conciseness
Figure 3: Comparison of 13 different data categories, including an evaluation of overall DDPs for the four requirements. Instagram achieves the highest overall DDP score for all the four requirements.
Important takeaways
[…]
  We found participants’ interpretation of the requirements are in line with that of European Data Protection Board. However, there seems to be conflicting interpretations of conciseness and transparency among both stakeholders.
[…]
5 RQ2(b): Improving DDPs comprehensibility
Figure 4: Comparison of watch history representation in our proposal and that of YouTube’s current DDP implementation.
(a) Clear and Plain Language
(b) Intelligible
(c) Transparency
(d) Conciseness
Figure 5: Percentage of votes for the proposal and the earlier winning platform across different categories for each requirement. Our proposal emerged as the top choice in all fields for the language and intelligible requirements.
In Section 4, we observed that none of the platforms adhere to the prescribed requirements of the GDPR properly across all data categories, highlighting the need for better representations. Motivated by this, here we utilize the different interpretations we elicit from survey participants to come up with novel recommendations that lead to better representations.
5.1 Methodology
To come up with an actionable recommendation for data representation, we take the following steps.
Step 1: Summarize the interpretations: As a first step, for each of the requirements, we collect the interpretations from our surveyed participants and that of the EDPB [14]. Next, we feed this entire set of 401 interpretations to a large language model (LLM) – Gemini 2.0 Flash [10]– using the Google AI Studio platform. We ask the LLM to generate a set of recommendations that encapsulate the most prevalent interpretations by summarizing the above responses. For reproducibility purposes, we make the prompts we used and the responses in  Appendix B. Note that in this approach, we include the interpretations of not only the data subjects, but also of data protection authorities (through the interpretation of EDPB).
Step 2: Generate recommended data field for each data category: In the second step, by providing the top recommendations for each of the requirement, we ask the LLM to make some assumptions. Specifically, we state that we are data engineers working in a short-format video platform operating in the European Union. Moreover, an end user requests us to provide them with their personal data for a given data category. Then, we ask the LLM to let us know of the specific data fields that we should provide to the end user. Subsequently, the LLM provides us with a set of recommendations for data fields of information in the queried category in the form of a JSON object (available in Appendix B).
Step 3: Manual inspection by researchers: In the final stage, three researchers who are part of this study and are conversant with DDPs of a whole array of platforms assess these recommendations. The primary goal of their assessment is to remove any instance of hallucination in the form of potentially irrelevant data fields that may have been recommended by the LLM. Next, given their expertise in understanding data donations in other platforms like Netflix, Prime video etc., we ask them to suggest any additions or removals to the recommended fields. By taking these steps, we ultimately come up with the final proposal of recommendations. Note that while Step 1 introduces the end users’ and data protection authorities’ interpretations, this step also instils the common interpretations of tech experts on other platforms and researchers who want to unravel the dynamics of platforms using these DDPs.
An example of the recommended representation for watch history and that of the most adhering representation among the existing one’s (YouTube’s) is shown in Figure 4. Note that apart from the details about the video watched, the proposal also shares the duration of the video and what fraction of it was watched by the user and on which device and network.
5.2 Evaluation of the proposed representation
To evaluate our proposed representation, we conducted another survey among 200 participants (who have not participated in the earlier survey mentioned in Section 4) from Germany, France, Spain, and Italy. We follow the identical participant recruitment strategy as elaborated in Section 4. Table 8 (See Appendix D) shows the user demographics. During the survey, we showed the participants two representations for the evaluated categories: (a) proposed representation: which we generate from the steps mentioned in Section 5.1, (b) winner representation: for each data category, the representation among Instagram/TikTok/YouTube which won the first preference along more requirements (i.e., the representation which participants suggested to be more adhering to the GDPR). We anonymize and randomly present each pair of representations to the survey participants. Then, we ask the participants to select one of the two representations that they find to be more intelligible/transparent/concise/in clear and plain language.
Observations We compare the representations of the data category fields of our proposal and that of the best among the existing short-format video platforms for 13 different categories across four GDPR requirements. [^5] Notice that the data categories are slightly different than the evaluations done in Section 4.3. We are not considering personal information, as there were no suggestions for adding or removing any fields. Instead, we include activity summary, which is provided only by TikTok, for comparison.  Figure 5 shows the percentage of votes for the two representations obtained in the 52 evaluations (13 data categories $\times$ 4 requirements), that we surveyed. We observe that the participants preferred our proposed representation in as many as 44 categories. However, their preferences vary for different requirements.
Language and intelligibility: For clarity and simplicity of language and intelligibility of the representation, our proposal gets upwards of 67% of the votes in most of the categories (except language in comments and intelligibility in autofill categories) suggesting our proposed representation is both intelligible and in clear and plain language.
Transparency: Our proposal is voted to be more transparent than the best of the existing short-format video platforms in 11 out of the 13 categories. Apart from search history, in all the other winning 10 categories, our representation gets upwards of 75% of the votes. In watch history, the split between our proposal and that of YouTube’s DDP is 47% and 53%, respectively, indicating that both representations are of comparable quality. However, in autofill information category, Instagram’s data representation gets nearly 93% of the votes from our survey participants. However, we argue that having less autofill information is indicative of less privacy-intrusive data collection. While our proposed recommendation includes only the email address, name, address, and phone number, that of Instagram contains much more sensitive information.
Conciseness: Notice that in Section 4.2, we mentioned conciseness and transparency, and to some extent intelligibility, are mutually competing requirements. We observe the same conflict in the results of this survey. Unsurprisingly, our proposal was preferred to be the most concise representation in merely 7 out of the 13 categories. Even then, in watch history, search history, login history, and device details – our proposed representation was found to be more concise by upwards of 75% of the participants. It is also worth mentioning that, much like the survey presented in Section 4.3, the participants from Germany have different preferences compared to those from other countries. Except for autofill information, over 66% of German participants preferred our proposed representation in all other categories.
Despite our proposed recommendation coming out as a significant improvement over any of the existing data representations, achieving this was not the main objective of this section. Our primary intention of this section has been to convey how, relatively easily, one can come up with more effective data representations by taking into account the interpretations of different stakeholders. This potentially indicates that currently, platforms are probably operating in silos without much deliberation and sincere effort into implementing Article 15 of the GDPR. At the same time, because of the mutually conflicting nature of the requirements, it is harder than anticipated to satisfy all the requirements with a single representation.
Important takeaways
[…]
  Taking into account the interpretations of different stakeholders (e.g., end users, legal scholars, technical experts, and researchers) can offer a significant improvement to the data representations across different categories.
[…]
6 RQ3: How reliable are the DDPs?
In this section, we investigate how reliable the DDPs are. We first define the requirements for quantifying the reliability of DDPs. Next, we outline the data collection employed to evaluate their reliability. Finally, we conduct a comprehensive assessment of the reliability of DDPs for the three platforms.
6.1 Requirements for reliability
86% of researchers who participated in a Data Donation Symposium reported encountering problems with the DDPs provided by various platforms [34]. The primary concerns identified include: (1) incompleteness, such as missing data types or limited time ranges, and (2) inaccuracy, such as wrong timestamps, incorrect ordering which renders the data unreliable to both end users and researchers. Therefore, due to the popularity of using DDPs to conduct research, there is a pressing need to systematically assess how reliable DDPs are. Hence we evaluate the reliability of DDPs on three requirements.
Completeness: Evaluates whether the platforms provide a comprehensive record of all user activities or whether the data is subject to sampling or omissions.
Correctness: Determines whether the sequence of recorded activities in the DDP accurately reflects the chronological order of events and whether any wrong entries are present.
Consistency: Assesses the uniformity of data across users and within users across different retrieval instances.
6.2 Experimental setup
We collect two types of data: (a) from real-world users, and (b) utilizing bot accounts.
Real-world users Inspired by prior work, we designed our own data donation website, which is privacy preserving and allowed us to obtain DDPs from real-world users [38]. To ensure user privacy, all collected data was anonymized and any PII was discarded on the front-end of our data donation website. During the data collection process, we intentionally mandated the donation of specific data categories for each platform. For TikTok and YouTube, video browsing history was mandatory. In the case of Instagram, the mandatory categories included (a) like history and (b) ads and topics, which consist of posts, ads, videos viewed, and ads clicked. Other data categories, such as search history, comments, etc., were optional and users provided them at their discretion. Private messages and other sensitive data were not collected. Users were compensated based on the categories and amount of data they donated. The maximum payment offered was \$16 for TikTok, \$9 for YouTube, and \$14 for Instagram data donations. Data collection for TikTok and YouTube took place between August and September 2024, while for Instagram, it was conducted between December 2024 and January 2025. We collected data from the same set of 33 users across all three platforms. 17 (51.5%) male and 16 (48.5%) female participants donated their data.
Bot accounts We automated the process of video browsing and video liking activities using pyautogui [2] to control mouse and keyboard and opencv-python [1] to detect the like button on the desktop screen. This approach differs from traditional browser automation [7], as we directly controlled the desktop environment instead of interacting with the browser via scripts. We captured interactions with the browser by downloading the HAR (HTTP Archive) files (HAR is a JSON-formatted file that logs all network traffic between the browser and the server). We viewed between 20 to 25 short-format videos in each video browsing session on each of the three platforms. Our bot accounts watched each video for a random duration, ranging from 15 to 60 seconds, and they randomly liked some of the videos. This approach allowed us to gather accurate ground truth data regarding user activities. The process was conducted for nearly thirty video browsing sessions between October 2024 and December 2024 for each platform. To conduct our analysis, we collected multiple DDP snapshots from each of these bot accounts.
6.3 Observations
Completeness: As mentioned in Section 3.2, Instagram and TikTok do not provide the entire watch history of a user, since they register on the platform, without any explanations. Hence, these DDPs are inherently incomplete and potentially violate Article 15(1)(d). However, with whatever data they share, to evaluate completeness, we utilize the data collected from the bot accounts. To this end, we compared the number of activities recorded in the HAR dump with those in the DDP for each video browsing session. We used the DDP that was closest in time to the corresponding HAR dump. Instagram offers two video browsing feeds: (1) home feed, and (2) reels feed. Interestingly, we found that videos watched under the reels feed were not present in the DDP, although like activities associated with them were recorded. For Instagram’s home feed, the percentage of recorded entries in the DDP for browsing videos and likes is 100%. On TikTok, all browsing and like activities were fully recorded, with a completeness rate of 100%. For YouTube, the recorded entries accounted for 99.5% of the activities observed in the HAR dump.
Correctness: After assessing the completeness of DDPs, we evaluated its correctness using the same data from bot accounts, by checking whether the order of contexts was maintained and whether any arbitrary entries were present in the DDP or not. To achieve this, we measured the Jaccard similarity for the following aspects between HAR and DDP for each activity type.
$\bullet$ Date: Consistency of recorded timestamps. We compared them up to the minute level to allow for slight variations.
$\bullet$ Context: Consistency of video IDs or author IDs.
$\bullet$ Overall: A combined evaluation of date and context. Table 4 presents the average Jaccard scores across the video browsing sessions for the three platforms. For Instagram, there is a drop in the Jaccard scores of browsing, because of changes in usernames, i.e, account username change or account deletion. For TikTok, Jaccard scores achieved 100% accuracy across both browsing and like activities. For YouTube, the average Jaccard scores are 100% for date, 99.83% for context, and 99.5% overall. The slight drop in the overall score is because of the 0.5% missing entries discussed earlier. For TikTok and YouTube, the recorded timestamps showed a minimal difference of $\pm$ 5 seconds compared to the HAR dump values for the same activity. However, for Instagram, the difference was as high as up to one minute, leading to a reduction in the accuracy for date and overall.
Platform (Activity) Date Context Overall
Instagram (Likes) 100% 100% 100%
Instagram (Browse) 96% 97% 91%
TikTok (Likes and Browse) 100% 100% 100%
YouTube (Browse) 100% 99.83% 99.5%
Table 4: Average Jaccard scores across different video browsing sessions for the three platforms. Concerning correctness, TikTok sits at the top, followed by YouTube and Instagram.
(a) Instagram
(b) TikTok
(c) YouTube
Figure 6: Ratio of entries retained across multiple snapshots for the same account: Instagram (like history), TikTok (video browsing history and like activities), and YouTube (video browsing history).
Consistency (across snapshots): Next, we determine whether a particular account receives consistent data when making multiple requests from the same platform. Towards this goal, we analyzed differences between multiple DDPs obtained from bot accounts. We examined whether the entries present in an earlier snapshot were retained in a later snapshot. For Instagram and TikTok, the snapshots are collected with a gap of 1 week. For YouTube, we use a gap of one month.
Figure 6 presents the ratio of overlapping entries between two DDPs to the total number of entries in the earlier DDP for three key aspects, i.e., date, context, and overall. For Instagram, despite using the same bot accounts, we observed variations in the covered duration of the video browsing history across multiple snapshots, ranging from 6 to 13 days. Hence, due to limited available video browsing history, we solely focus on analyzing the like history.  Figure 6(a) shows the results for Instagram. We identify some missing entries in the later snapshots. Specifically, 6% of the entries are found to be missing from snapshot 3 to snapshot 4.  Figure 6(b) illustrate the results for video browsing history and like activities on TikTok. Based on our findings, we observe that TikTok provides consistent data across multiple retrieval attempts. Lastly, in  Figure 6(c), we observe that approximately 12% of entries were missing for snapshot pairs (1, 3) for YouTube. Among these, nearly 62% entries were advertisements, while the other videos got deleted from the platform itself.
Consistency (across users): To assess whether all users receive a similar amount of data, we analyzed video browsing history, search history, and like history obtained from real-world users across the three platforms. First, we measured the duration of data that was contained in each activity history by calculating the time difference between the earliest and the latest entries in the respective activity history lists. For Instagram, we defined browsing history as the combination of ads viewed, posts viewed, and videos watched, while we included both keyword or phrase searches and user searches in the search history. To understand the extent to which the distribution of provided data to the users differs, we analyzed the cumulative distribution function (CDF) of the durations of video browsing, search, and like history across the platforms.
Figure 7 presents a comparative view of activity durations across users. For Instagram, we observe distinct clusters of users with browsing history durations concentrated around 6 and 13 days (Figure 7(a)). This non-uniform distribution indicates that some Instagram end users receive significantly different amounts of data. Similarly, on TikTok (Figure 7(c)), video browsing and search history durations exhibit substantial variation. Specifically, some end-users had data spanning approximately 180 days, while others had nearly 450 days of recorded video browsing and search activities. To ensure a fair comparison, we specifically restricted our analysis to TikTok users whose earliest recorded activity occurred before March 2023. This criterion allows us to focus on the subset of users who have had ample time to accumulate browsing and search history. As a result, we analyzed data from 23 TikTok users. These findings highlight TikTok’s tendency to provide data inconsistently among its users. However, such kind of clusters were not observed for like history on both Instagram and TikTok. The analysis of video browsing and search history on YouTube (Figure 7(d)) does not reveal any discernible patterns, suggesting that the platform may provide DDPs more consistently to its end users.
(a) Instagram - Browse, Search
(b) Instagram - Likes
(c) TikTok
(d) YouTube
Figure 7: Plots illustrating the CDF of data duration provided to users for three activities – search, browse, and like history. On Instagram, clusters in browsing history are observed at 6 and 13 days, while on TikTok, clusters appear at 180 and 455 days for both search and browse history.
Important takeaways
[…]
  In terms of correctness and completeness, TikTok provides the most reliable DDPs. Also, its DDPs are consistent across multiple retrievals.
[…]
7 Concluding discussion
Summary of the insights: To the best of our knowledge, ours is the first systematic attempt to evaluate the implementation of the right of access (Article 15) across platforms offering similar services. Our detailed evaluation of Instagram, TikTok, and YouTube, focusing on their DDPs, reveals that Instagram and TikTok share more data categories, while YouTube’s shared data is limited. The only underlying similarity that cuts across platforms is that of non-disclosing the purpose of data collection and processing. While Instagram’s data representation is deemed most comprehensible, TikTok’s is found to be the most reliable. Finally, a proposed data representation, leveraging interpretations of different important stakeholders, is found to be preferred by participants across multiple countries along the different requirements.
Limitations of the current work: Like any other work it has its own share of limitations. The survey population could have been more diverse and representative of common users on the street. However, note that our work is comparative by nature and we do not expect that the choice of participants will impact the relative rankings and comparisons presented across the paper. Further, our study solely focuses on data shared by platforms upon request and not the data that is actually collected from the users. Including this information may further enrich the quality of the work.
Recommendations for stakeholders: Platforms should improve the consistency among different categories of data that they share with end-users. The way different related categories of data are being shared differently (e.g., like vs. browsing history on Instagram) raises concerns about the lack of coordination within the platform while implementing the regulations. The other significant room for improvement for platforms is increasing the comprehensibility of the shared data. As we show in our work if we consider the interpretation of different stakeholders, the resulting representations can already be a significant improvement on the current representations.
The current differential and ineffective implementation not only affects end-users’ rights of access but also affects their rights to port their data (Article 20). Such rights of users will only be useful if data protection authorities provide more detailed technical specifications and/or standardization for implementation. As Section 5 shows involving different stakeholders may further improve the standard of implementations. Moreover, authorities need to enforce these requirements–otherwise platforms may not have any ramifications to worry about if they do not comply with them. To this end, the current study also indicates how effective enforcement of the GDPR rights may benefit from computer science research by relying on audit strategies to observe compliance at scale.
Acknowledgments
We would like to extend our sincere gratitude to Prof. Christoph Engel (Max Planck Institute for research on Collective Goods) and Prof. Cristoph Sorge (University of Saarland) for their valuable feedback during the course of this work. Additionally, Ingmar Weber is supported by funding from the Alexander von Humboldt Foundation and its founder, the Federal Ministry of Education and Research (Bundesministerium für Bildung und Forschung).
8 Ethics considerations
We placed the highest priority on ensuring that our research project adheres to standard ethical guidelines. To this end, and before initiating our study, we acquired approval from the Ethical Review Board (ERB) of our University. Before surveying a participant or obtaining their DDPs, we provided them with an online consent form and asked them to meticulously review the outlined terms in the consent form and explicitly provide us with their consent to these terms. We refrained from collecting any PII from our participants who took part in our surveys and donated us their DDPs. Specifically, we removed any available PII inside the DDPs at the front-end and also anonymize them before they are sent to our back-end server. Furthermore, according to the stated terms and conditions in the ERB approval, we took stringent measures to protect the privacy of our participants. In particular, we would not share their donated anonymized DDPs with any third parties, and we would completely remove these DDPs within a specified time after the conclusion of our study.
When conducting our surveys, we ensured that the questions did not pose any potential risks to our participants. Specifically, when surveying our participants on different representations of the DDPs, we filled out data categories based on sample DDPs that our research team obtained from their own personal accounts on the studied platforms. Moreover, we checked the contents of these data categories and verified that they do not endanger our participants in any way. We utilized bot accounts that exclusively interacted with publicly available content on the platforms to carry out one part of our experiments in Section 6. We took multiple steps to minimize the inadvertent consequences of running these bot accounts. To this extent, we intentionally distributed the incurred overhead of our experiment over multiple short browsing sessions to prevent causing any sort of disruptions to the short-format video platforms. Specifically, on a single day of our experiment, a bot account had only one video browsing session and viewed up to a maximum of 25 short-format videos. Moreover, although our bot accounts tangibly engaged with some videos by liking them, we used only three bot accounts across all the platforms, so we do not foresee any potential harm resulting from these activities. Furthermore, although the Terms of Service of the studied platforms prohibit the usage of automated bots, the benefits of our study outweigh any potential risks associated with the activities conducted by our bot accounts. Finally, we affirm that we fully conformed to ethical research standards throughout our work [28]. This includes but is not limited to, presenting our findings in aggregate form and guaranteeing the anonymity of all our participants.
9 Open science
We comply with standard open science guidelines to the fullest extent permitted by the terms and conditions outlined in the ERB approval that we obtained. To this end, we publicly share multiple materials that we utilized to carry out our research project in References. Specifically, we provide details on the surveys which we conducted in Appendix B. These details include the exact questions and choices that we gave to our participants. Moreover, we present details on the prompts we fed to the LLM and the responses that we got from the LLM also in Appendix B.
As previously stated in Section 8, according to the terms outlined in our approved ERB proposal, we would remove all the DDPs that we collected in Section 6 from our database after the completion of our research project. Therefore, we cannot make these anonymized DDPs publicly available.
Appendix A Content of DDPs provided by YouTube, TikTok, and Instagram.
Table 5 presents an overview of the content included in the DDPs offered by these three platforms. Note that although Instagram has cookies field it just shows the last four digits of the cookie name and no other details.
Information Type TikTok Instagram YouTube Minimum expected fields
User’s usage Watch Y N Y Content Id, ts
Search Y Y Y Query term, ts
Comment Y Y Y Comment text, Content Id, ts
Like Y Y $Y^{g}$ Content Id, ts
Messages Y Y NA Content, User Id, ts
Save/Favourite Y Y Y Content Id, ts
Share(In-app) Y Y NA Content Id, ts
Share(Across-app) Y N N Content Id, ts
Interests/Topics Y Y N List of topics
Time spent $N^{*}$ $N^{*}$ $N^{*}$ Duration/frequency
User’s content Media Y Y Y Media file/URL
Text details Y Y Y title
Location Y Y Y Some place identifiers
Date time Y Y Y ts
Device - Y N Device model, OS
Other user interactions $N^{*}$ $N^{*}$ $N^{*}$ Likes, Comments
Personal details Account details Y Y $Y^{g}$ Username, DOB, Email, Profile photo
Connections Y Y Y Username, ts
Login history Y Y $Y^{g}$ IP, ts
Current devices Y Y $Y^{g}$ User agent
Current camera NA Y NA Version/type
Location Y Y $Y^{g}$ Place identifiers
Account changes - Y N Type, Old, New values, ts
Advertisements Ads viewed N N Y Content Id, ts
Personlization ${N^{*}}$ N N Reasons why the ad was shown
Access to your data N Y N Which and how (in store visit etc.)
Miscellaneous Off-platform Y Y N Platform, ts, activity
Link history - Y - Link, ts
Cookies - Y - -
Table 5: Data Transparency: Overview of collected and shared information. ${N^{*}}$ denotes the details are found in the app but not in the GDPR dump. ${Y^{g}}$ denotes the details are found in google’s DDP, but not in YouTube’s DDP. (ts-TimeStamp)
Appendix B Survey questions, prompts, and responses from an LLM.
We included the following in the supplementary material [^6] https://zenodo.org/records/14721885 :
1. Survey questions used for Section 4.
2. Overall DDPs used for evaluation in Section 4.
2. Prompts provided to LLM and responses from LLM used in Section 5.
4. Survey question used for Section 5.
Appendix C Comprehensibility of DDPs
Table 6 presents the distribution of the participants recruited for the survey in Section 4.
Attribute Type Count Percentage
Gender Male 228 57.0
Female 161 40.3
Other 9 2.3
Prefer not to say 2 0.5
Age 18-25 95 23.7
25-30 104 26.0
30-40 114 28.5
40+ 87 21.8
Country France 100 25.0
Germany 100 25.0
Spain 100 25.0
Italy 100 25.0
Table 6: Distribution of participants in survey reported in Section 4 based on their self-reported gender, age, and country of residence.
C.1 Evaluation for the requirements across the platforms
Table 7 shows the data categories considered for evaluating the comprehensibility of DDPs.
Category What does it mean?
Watch history List of contents watched by a user on the platform.
Search history List of terms/users searched by a user on the platform.
Comments List of comments written by a user on others posts/videos on the platform.
Saved List of contents saved by a user for future reference on the platform.
Connections List of other users connected to a user on the platform.
Location The most recent location as recorded by the platform.
Autofill information Personally Identifiable Information including name, mail id, phone number, address etc. often collected implicitly.
Off-platform activity User’s activity data shared by other platforms with the platforms under consideration.
Devices Details of the devices that a user has used to connect to the platform.
Login history Device and network from which the user has logged into the platforms.
Like history List of contents liked by a user.
Personal information Personally Identifiable Information, including name, mail ID, phone number, address, etc., shared by the user.
User’s content List of contents uploaded or posted by the user on the platform.
Table 7: A list of categories of data that were considered in the adherence evaluation phase of the user survey.
Figure 8 shows the representation of watch history for the three platforms and sample question format.
Figure 8: Figure representing how the watch history from the three platforms was displayed to users for comparison and to evaluate the four properties: conciseness, clear and plain language, intelligibility, and transparency.
Appendix D Improved DDPs comprehensibility
Table 8 presents the distribution of the participants recruited for the survey in Section 5.
Attribute Type Count Percentage
Gender Male 100 50.0
Female 99 49.5
Prefer not to say 1 0.5
Age 18-25 54 27.0
25-30 57 28.5
30-40 57 28.5
40+ 32 16.0
Country France 50 25.0
Germany 50 25.0
Spain 50 25.0
Italy 50 25.0
Table 8: Participant distribution in the survey reported in Section 5.2 based on their self-reported gender, age, and country.

### v2

@@
 Abstract
-The comprehensibility and reliability of data download packages (DDPs) provided under the General Data Protection Regulation’s (GDPR) right of access are vital for both individuals and researchers.
-These DDPs enable users to understand and control their personal data, yet issues like complexity and incomplete information often limit their utility.
-Also, despite their growing use in research to study emerging online phenomena, little attention has been given to systematically assessing the reliability and comprehensibility of DDPs.
-To bridge this research gap, in this work, we perform a comparative analysis to assess the comprehensibility and reliability of DDPs provided by three major social media platforms, namely, TikTok, Instagram, and YouTube.
-By recruiting 400 participants across four countries, we assess the comprehensibility of DDPs across various requirements, including conciseness, transparency, intelligibility, and clear and plain language.
-Also, by leveraging automated bots and user-donated DDPs, we evaluate the reliability of DDPs across the three platforms.
-Among other things, we find notable differences across the three platforms in the data categories included in DDPs, inconsistencies in adherence to the GDPR requirements, and gaps in the reliability of the DDPs across platforms.
-Finally, using large language models, we demonstrate the feasibility of easily providing more comprehensible DDPs.
-  *These authors contributed equally to this work.
+The GDPR’s Right of Access aims to empower users with control over their personal data via Data Download Packages (DDPs).
+However, their effectiveness is often compromised by inconsistent platform implementations, questionable data reliability, and poor user comprehensibility.
+This paper conducts a comprehensive audit of DDPs from three social media platforms (TikTok, Instagram, and YouTube) to systematically assess these critical drawbacks.
+Despite offering similar services, we find that these platforms demonstrate significant inconsistencies in implementing the Right of Access, evident in varying levels of shared data.
+Critically, the failure to disclose processing purposes, retention periods, and other third-party data recipients serves as a further indicator of non-compliance.
+Our reliability evaluations, using bots and user-donated data, reveal that while TikTok’s DDPs offer more consistent and complete data, others exhibit notable shortcomings.
+Similarly, our assessment of comprehensibility, based on surveys with 400 participants, indicates that current DDPs substantially fall short of GDPR’s standards.
+To improve the comprehensibility, we propose and demonstrate a two-layered approach by: (1) enhancing the data representation itself using stakeholder interpretations; and (2) incorporating a user-friendly extension (Know Your Data) for intuitive data visualization where users can control the level of transparency they prefer.
+Our findings underscore the need for clearer and non-conflicting regulatory guidance, stricter enforcement, and platform commitment to realize the goal of GDPR’s Right of Access.
+The supplementary materials and code are available at https://github.com/saikeerthana00/GDPR_Compliance. This work has been accepted at IEEE S&P 2026 as a full paper.
 1 Introduction
-Ensuring compliance with the General Data Protection Regulation (GDPR) [12] is essential for platforms handling personal data, as it establishes legal requirements for data processing, security, and user rights.
-One of these rights is the right of access by data subjects, which enlists provisions whereby data subjects (i.e., end users) can request access to their data being processed by platforms (Article 15(3) GDPR).
-Platforms have to comply with these requests within a stipulated period of one month and provide the user with a copy of their data.
-Usually, platforms provide this data in the form of data download packages (DDPs) [34].
-These DDPs often include information about the users’ activities on and off the platform, device and app information, inferred preferences, and other information depending on the business of the platform.
-Implementation across platforms: In fact, the GDPR stipulates that information shall be provided to users in commonly used electronic forms (Art.
-15(3)(2) GDPR).
-Such an implementation is crucial for enabling users to exercise their right to data portability (Article 20), which allows end users to transmit data from one platform to the other without hindrance [12].
-However, the GDPR does not provide explicit standards for the structure or content of DDPs, leading to varying implementations of Article 15 across platforms [8].
-Understandably, platforms that offer different types of services interpret and implement the GDPR provisions differently.
+The General Data Protection Regulation (GDPR) [10] establishes legal requirements for data processing, security, and user rights.
+One of these rights is the Right of Access by data subjects, which includes provisions that allow data subjects (i.e., users) to request access to their data being processed by platforms (Article 15(3) GDPR).
+Along with a copy of such data, GDPR also requires platforms to communicate the purpose of data collection, period of retention, and other recipients [^1] As per Article 4 (9) of the GDPR, ‘recipient’ means a natural or legal person, public authority, agency or body, to which personal data are disclosed [10].
+of such data (Article 15(1) GDPR).
+Platforms have to comply with these requests within a stipulated period of one month.
+Platforms usually provide this data in the form of data download packages (DDPs) [33].
+Implementation across similar platforms: However, the GDPR does not provide standards for the structure or content of DDPs, leading to varying implementations of Article 15 across platforms [6].
+Understandably, platforms that offer different types of services are expected to interpret and implement these GDPR provisions differently.
 For instance, the DDP of a digital marketplace like Amazon is naturally distinct from that of a digital public space like TikTok.
 However, platforms offering similar services should be able to provide comparable information to users.
-For example, platforms providing short-format video streaming services should, in principle, offer similar details through their DDPs.
-To the best of our knowledge, such a comparative study of the implementation of the right of access under the GDPR is rare in the literature.
-This observation leads to a key unexplored research question as follows: RQ1: How similarly do platforms offering similar services implement the right of access under the GDPR (Article 15(3))?
-Desideratum 1: Comprehensibility: Irrespective of the answer to RQ1, these implementations (DDPs) should adhere to a set of requirements outlined in the GDPR to ensure user comprehensibility.
-According to Article 12(1), these DDPs should be (a) concise, (b) transparent, (c) intelligible, (d) easily accessible, and (e) using clear and plain language [12].
-However, neither the GDPR nor accompanying guidelines [14] provide a clear technical definition of these criteria.
-Further, adherence to these requirements in the current implementations also remains unexplored in the research community.
+For example, platforms providing short-format video streaming services should, in principle, offer similar details in their DDPs.
+Such consistent implementation across similar platforms is one of the important factors for enabling users to exercise their Right to Data Portability (Article 20 GDPR), which allows users to transmit their data across platforms without hindrance [10].
+Therefore, there is a need for systematic and comparative audit of implementations across platforms.
+To the best of our knowledge, such a comparative study of the implementation of the GDPR’s Right of Access across similar platforms is rare in the literature.
+These concerns lead to the following key unexplored research question:
+RQ1: How similarly do platforms offering similar services implement and thus comply with GDPR’s Right of Access?
+Reliability of DDPs’ content: Irrespective of the similarity of implementations, a fundamental concern is the reliability of the content in DDPs.
+Lack of reliability of the DDPs undermines the overarching purpose of the Right of Access and severely compromises their utility towards users.
+Furthermore, unreliable DDPs can create a distorted understanding of data processed by platforms, limiting users’ digital autonomy and right to digital privacy.
+Such lack of reliability can also have ramifications for other stakeholders too.
+Researchers increasingly rely on these DDPs for studying the dynamic interplay between population and platforms [39, 38, 35, 36].
+To this end, incorrect or inconsistent information in platforms’ DDPs risks the insights drawn from such studies being fundamentally flawed and misleading.
+In fact, the above concern is no longer a mere theoretical one.
+A recently conducted survey among researchers who have used DDP in their research indicates that 86% of them have reliability concerns with DDPs [33].
+Their concerns are primarily centered around the following dimensions – (a) completeness: ensuring all relevant data is present in DDPs for the stipulated period, (b) correctness: verifying the accuracy of the information, such as timestamps and event sequences recorded in DDPs, and (c) consistency: assessing uniformity of DDPs across different requests for the same user or across users in different locations.
+While researchers have reported these concerns, there is no prior systematic audit of the reliability of the DDPs, leading to our second research question:
+RQ2: How reliable is the information within the DDPs provided by platforms?
+Comprehensibility of DDPs: Even if DDPs are to contain accurate and complete information, their utility would be severely limited if users cannot easily understand the content presented to them i.e., if DDPs are not comprehensible for end-users.
+In fact, to ensure comprehensibility of the communication, Article 12(1) of the GDPR states these DDPs should be (a) concise, (b) transparent, (c) intelligible, (d) easily accessible, and (e) using clear and plain language [10].
+However, neither the GDPR nor accompanying guidelines [12] provide any clear technical definition of these requirements.
+Additionally, adherence to these requirements in the current implementations also remains unexplored in the research community.
 The authors in [6] came closest to understanding the question of adherence, but they restricted their analyses to conciseness.
-This gap leads us to the second research question in this paper: RQ2: How comprehensible are the DDPs provided by platforms under Articles 12 and 15 of the GDPR?
-Desideratum 2: Reliability: Finally, another important expectation for the effective implementation of the provision is that the information in the DDPs is reliable.
-With the increased restriction on access to platform data for research [33, 4], many researchers are relying on data elicited through exercising the right of access for understanding the dynamic interplay between population and platforms.
-In these so-called “data donations” based studies, consenting participants first request their own data from the platforms, and then provide this data to researchers, typically for financial compensation.
-Several recent scientific studies use this setup to explore fundamental questions about digital platforms— such as user engagement [38], content recommendation [35], and ad targeting processes [36] — and even questions beyond online activities, e.g., offline risks [3] or the likelihood of addiction [37] using the resultant DDPs.
-However, if the reliability of the information provided by the platforms is compromised, then the insights drawn from such data sources might be invalid.
-To this end, a recently conducted survey among researchers indicates that 86% of researchers report concerns regarding the reliability of the data collected using DDPs [34].
-The most important dimensions of reliability that researchers are concerned about include: (a) completeness, (b) correctness, and (c) consistency of the shared information.
-This brings us to our final research question: RQ3: How reliable is the information within the DDPs provided by platforms under Article 15(3) of the GDPR?
-The current work: Although the current work could have been conducted on any digital platform, due to the growing popularity in consumption of short–format videos [38, 25] and the impact they have on the population [13], we investigate the aforementioned research questions on Instagram, TikTok, and YouTube.
+This gap leads to the third research question in this paper:
+RQ3: (a) How comprehensible are the DDPs provided by platforms and (b) how can their comprehensibility be improved?
+Figure 1: Pipeline to evaluate comprehensibility and reliability of the implementation of Article 15(3) of the GDPR.
+The current work: Although the current work could have been conducted on any digital platform, due to the growing popularity of short–format videos [39, 23] and their impact [11], we investigate the discussed research questions on TikTok, Instagram, and YouTube.
 Figure 1 describes the pipeline of our conducted research.
-After analyzing the DDPs shared by these platforms, we conducted an extensive survey among 400 participants from Germany, France, Italy, and Spain to evaluate their assessment of the comprehensibility of the respective DDPs.
-Further, by utilizing (a) participants’ and (b) European Data Protection Board’s interpretations of the requirements, (c) current advancements in large language models (LLMs), and (d) our technical expertise, we propose data representations for the different categories of information.
-These recommended representations are then evaluated against the best representations of respective categories in the current implementation of the platforms by 200 participants.
-To answer the questions regarding the correctness and completeness of the DDPs, we browse the three platforms using sock-puppet accounts to (a) automatically log the browsing behavior and (b) request the DDPs from each of the platforms.
+To audit reliability of DDPs, we browse the three platforms using sock-puppet accounts to (a) automatically browse and log the behavior and (b) request the corresponding DDPs from each of the platforms.
+We quantify correctness and completeness by comparing our log with the DDPs from each platform.
 We repeat this process for a period of one month to understand the consistency in the shared DDPs for the accounts.
-In addition, to better understand the consistency across various accounts, we also collect DDPs requested by real-world participants from the three platforms.
-Figure 1: Our pipeline to evaluate comprehensibility and reliability of the current evaluation of Article 15(3) of the GDPR.
-We summarize our major findings with respect to each research question below.
-RQ1: The studied platforms implement Article 15 of the GDPR differently: Instagram, TikTok, and Youtube share different categories of data in their DDPs.
-While Instagram and TikTok share more data with their users, YouTube’s shared data is rather limited.
-Irrespective of the amount of data, the only underlying similarity is that none of the platforms share the purpose of data collection and processing in their DDPs.
-RQ2: Instagram’s current data representation is the most comprehensible: Based on the preference votes of 400 participants, we find Instagram’s DDP adheres to GDPR’s requirements more than the other two platforms.
-RQ2: Considering interpretation from different stakeholders significantly improves the data representation: 200 survey participants prefer our proposed data representation to current representations along all requirements except conciseness.
-In fact, our proposal is preferred in as many as 44 out of the 52 evaluations (13 categories $\times$ 4 requirements).
-RQ3: TikTok’s DDP is the most reliable as per our evaluation: Our analyses using sock-puppet accounts reveal TikTok’s DDP presents a complete, correct, and consistent representation of our browsing behavior.
-Instagram and YouTube exhibit varying degrees of missing data across DDPs collected over the period of our study.
-However, the analyses among real users show Instagram and TikTok share data for different periods in their DDPs for different categories, whereas YouTube shares all the data types for similar duration.
-To the best of our knowledge, this study is the first detailed evaluation of adherence of different requirements mentioned in the GDPR.
-The insights drawn from our study further underline the need for standardization and follow-up on the implementations to ensure that the primary goals of data protection laws are achieved.
+Also, to better understand the consistency across various accounts, we collect DDPs requested by real-world participants from the three platforms.
+To assess the comprehensibility of DDPs, we conduct an extensive survey among 400 participants from Germany, France, Italy, and Spain.
+Beyond understanding the comprehensibility of the DDPs, we also ask participants about their interpretation of different GDPR prescribed requirements.
+One of the key observations in this survey is that some of the requirements are at odds with each other.
+To this end, by utilizing interpretations of participants and European Data Protection Board, we propose data representations for the different categories of information that aim at striking a better balance among these requirements.
+Furthermore, to improve the comprehensibility of the DDPs, we propose a browser extension (Know Your Data) which (a) enables users to request and download their data in one click; (b) provides them the autonomy to choose their desired level of transparency in data visualization.
+The code is made publicly available.
+Below, we summarize our major findings.
+RQ1: Instagram, TikTok, and YouTube interpret and implement GDPR’s Right of Access differently, with YouTube offering surprisingly limited data compared to the others.
+Critically, none adhere to the fundamental requirement of (a) disclosing the purpose of data collection, (b) duration of retention, (c) other entities with whom the data is shared for different data categories within DDPs.
+RQ2: Our analyses using sock-puppet accounts reveal TikTok’s DDP presents a complete, correct, and consistent representation of our logged browsing behavior.
+Instagram and YouTube exhibit varying degrees of missing data across DDPs collected during our study.
+However, the analyses among real users show Instagram and TikTok share data for different periods in their DDPs for different categories, whereas YouTube shares all data for a similar duration.
+RQ3(a): In our user survey, we find Instagram’s DDP to be more comprehensible than that of the others.
+However, survey participants’ interpretation of the different requirements reveals some of them may have conflicting goals; e.g., while participants interpret conciseness to prioritize relevant information, they interpret transparency to provide complete information with full disclosure.
+RQ3(b): Improving comprehensibility requires a layered approach.
+One can improve the data representation themselves by accounting for interpretations from different stakeholders to achieve a better trade-off between the conflicting requirements.
+Furthermore, to improve the understandability, an additional layer of visualization in the form of a dashboard may improve the current state-of-the-art by enabling users to decide the level of conciseness/transparency they prefer.
+Implications: Our findings reveal several instances of platform-specific and systemic non-compliance, providing regulators with starting points to investigate the platforms.
+Furthermore, insights drawn from our surveys highlight the conflicting nature of GDPR’s comprehensibility requirements making it difficult for platforms to adhere to all of them at once.
+At the same time, our recommendations underline the need for standardization and improved DDP communication through dashboards to ensure that the primary goals of data protection laws are achieved.
 2 Background and related work
-To position the current work in the perspective of the literature, we present a brief overview of the following: (a) GDPR and the rights of access, (b) leveraging data download packages for platform audits, (c) audits of implementations of provisions of the GDPR.
-GDPR and the right of access: The GDPR is a landmark privacy law designed to give individuals greater control over their personal data and to standardize data protection rules across the European Union (EU) [12].
-While the GDPR has many provisions and rights enlisted, our work focuses on Article 15(3): the right of access, which allows users to request a copy of their data collected by online platforms.
+GDPR and rights of data subjects: The GDPR is a landmark privacy law designed to give individuals greater control over their personal data and to standardize data protection rules across the European Union (EU) [10].
+While the GDPR has many provisions and rights enlisted, our work focuses on Article 15: the Right of Access, which allows users to access their data collected and processed by online platforms.
 Most platforms implement this requirement by providing a Data Download Package (DDP) upon request.
@@
 Specifically, Article 12(1) requires a DDP to be concise, intelligible, transparent, accessible, and in clear and plain language [10].
-Moreover, the GDPR gives the right to end-users to be able to port their data from one platform to another under Article 20.
+Moreover, Article 20 gives the right to end-users to port their data from one platform to another.
 However, if platforms offering similar services implement these rights in significantly different ways, platforms will essentially deprive end users of their right to portability.
-Leveraging data download packages: The richness of DDPs has enabled new research directions previously deemed infeasible.
-Indeed, many recent studies are utilizing them through donations made from participants with their explicit consent.
-To this extent, researchers have used DDPs to conduct research on topics related to personal health and safety [20, 37, 3], news and politics [6, 17, 18], auditing recommendation algorithms [24, 35], analyzing user behavior on social media platforms [38, 15], and understanding ad targeting types and user perceptions [36].
-Audits of implementations of GDPR: Many studies have investigated (a) whether consent forms are designed properly by platforms [29, 30, 23, 9], (b) whether platforms are undermining user privacy under the pretense of ‘legitimate interest’ [31, 21, 22], (c) inconsistencies in platform’s data withdrawal behavior [11].
-However, little attention has been given to platforms’ implementation of the right to access.
-A recent study shows that end users often view DDPs as overwhelming and unclear, leaving key questions unanswered while triggering privacy concerns [8].
-Moreover, researchers have expressed concerns about the reliability of the DDPs and have faced challenges in using them [34, 19].
-The current work: To the best of our knowledge, our work is one of the first studies that tries to unravel the current implementations of Article 15(3) of the GDPR from the lens of comprehensibility and reliability.
-Furthermore, our proposed recommendations for the potential implementation of DDPs may not only improve the current implementations, but also help in the implementation of other rights provided to users under the GDPR, such as the right of data portability.
-3 RQ1: How similarly do platforms offering similar services implement GDPR Article 15?
-Before analyzing the comprehensibility and reliability of DDPs, we first look into the content of the DDPs provided by each of the platforms (i.e., we compare how the platforms are currently implementing GDPR Article 15(3)).
+Leveraging DDPs for platform audits: Even at the current level of implementation, the richness of DDPs has enabled new research directions previously deemed infeasible.
+Many recent studies are utilizing them through donations made from participants with their explicit consent.
+To this extent, researchers have used DDPs to conduct research on topics related to personal health and safety [18, 38, 3], news and politics [5, 15, 16], auditing recommendation algorithms [22, 35], analyzing user behavior [39, 13], and ad targeting on social media platforms [36].
+GDPR compliance audits: Since the enforcement, numerous academic evaluations have been conducted to assess compliance with different provisions of GDPR.
+Prior works have investigated (a) effective implementation of data portability [37, 32], (b) ethical design of consent forms [28, 29, 21, 7], (c) privacy violations under the pretense of ‘legitimate interest’ [30, 19, 20], and (d) inconsistencies in platform’s data withdrawal behavior [9].
+Current work: However, little attention has been given to platforms’ implementation of the right to access.
+While there are some studies raising privacy [6] and reliability concerns [33, 17], none of the studies systematically audit the Right of Access on the legal requirements proposed by the GDPR itself.
+To the best of our knowledge, this work is one of the first studies that tries to empirically evaluate the current implementations of Article 15 of the GDPR from the combined lens of comprehensibility and reliability.
+Besides auditing the compliance, we also propose a number of recommendations for policymakers and platforms alike to improve the comprehensibility for end-users.
+3 Auditing compliance of similar platforms
+Before analyzing reliability and comprehensibility, we first compare how platforms are currently implementing GDPR Article 15 by looking at the content of their DDPs and asses their compliance with the prescribed requirements.
 3.1 Platforms under consideration
-Since the primary focus of the study is to understand how similar platforms implement the same provision in the GDPR, we decided to focus on multiple platforms providing similar services.
-To this end, there are many options for digital streaming platforms (e.g., Netflix, Disney, etc.), digital marketplaces (e.g., Amazon, Zalando, etc.), and social networking platforms (e.g., Facebook, X, LinkedIn, etc.).
 In this study, we focus on three platforms that provide short-format video streaming services: TikTok, Instagram, and YouTube.
-Our motivation to study these platforms stems from the recent increase in popularity of short-format video platforms [38, 25].
-Furthermore, all three platforms have been designated as Very Large Online Platforms (VLOPs) by the EU under the recently enacted Digital Services Act (DSA), indicating the far-reaching impact that these platforms have on the European society [13].
-TikTok, Instagram, and YouTube implement Article 15(3) of the GDPR, through which they share DDPs with their end-users upon request.
-In the current implementations, platforms share data primarily in two formats (a) machine-readable – which is usually a single JSON or a collection of JSON files; (b) human-readable – where Instagram and YouTube opt for HTML version, whereas TikTok shares it in TXT files segregated across many directories.
-If one takes a closer look at the DDPs received from any of the platforms, their contents can be broadly divided into the following categories: (a) user’s usage, (b) user’s content, (c) personal details, (d) advertisements, (e) miscellaneous.
-Next, we examine the DDPs that Instagram, TikTok, and YouTube provide and observe their content and presentation.
-Most of the observations discussed below are summarized in Table 5 (see Appendix A).
-3.2 Information regarding users’ usage
-One of the most important facets of the information contained within the DDPs is how a user uses the social media platform.
-Such usage information may include the content that a user watches, searches for, or engages with implicitly or explicitly.
-Watch history: Watch history is an ordered list of all the content that a user has watched before requesting the data.
-DDPs of all three platforms contain watch history data of the data subject.
-However, as Table 1 shows the data provided across the three platforms are widely varying.
-Among these, two points are strikingly different – (a) while YouTube provides this data for the entire lifetime of a user on the platform, Instagram provides data of at most the last two weeks and TikTok provides data for around six months before the data request (without any clear explanation for the reduced duration); (b) while TikTok and YouTube provide the watch history all at once, Instagram segregates it into three files based on the type of content, i.e., ads, posts, and videos watched by the user.
+Our motivation to study these platforms stems from the recent popularity of short-format video platforms [39, 23].
+Furthermore, all three platforms have been designated as Very Large Online Platforms (VLOPs) by the European Union (EU) under the recently enacted Digital Services Act (DSA), indicating the far-reaching impact that these platforms have on the European society [11].
+What do these platforms share?: Upon users’ request, these platforms share Data Download Packages (DDPs) primarily in two formats (a) machine-readable – a single or a collection of JSON/CSV file(s); (b) human-readable – Instagram and YouTube opt for HTML version, whereas TikTok shares in TXT files segregated across directories.
+If one takes a closer look at the DDPs received from any of the platforms, their contents can be broadly divided into the following categories: (a) user’s engagement, (b) personal details, (c) advertisements etc.
+Next, we examine the DDPs that TikTok, Instagram, and YouTube provide and observe their degree of compliance.
 Feature YouTube Instagram TikTok
@@
 Single file
-Table 1: Comparison of watch history data across platforms.
-Like history: Similar to watch history, like history is an ordered list of all the contents that a user has liked before requesting the data.
-The details across the three platforms are noted in Table 2.
-The table demonstrates that there is a clear lack of consistency in how platforms currently implement Article 15(3) GDPR.
-While one platform (YouTube) records the like history on its application, it does not share this information in the YouTube DDP  [^1] Although Like history is not there in YouTube ddp, it is included in the Google DDP..
-Another platform (Instagram) provides the watch history for just a couple of weeks without any link to the watched content, but it provides the like history of a lifetime, including references to the liked content.
+TABLE I: Comparison of watch history across platforms.
 Platform
 Provides
-like
-history?
 Details provided
@@
 version but does not share it in the DDP
-Table 2: Comparison of like history details across platforms.
-Time spent on platform: While all three platforms have a detailed dashboard that mentions the average time spent and the exact amount of time users spend on the application daily, the same information is not provided in the DDP in response to Article 15 GDPR.
-On a positive note, TikTok recently added a new field in the data, namely “Activity Summary,” which enlists how many videos a user has commented on, shared, or watched until the end since their registration.
-Other usage activities on platform:  Table 5 (in Appendix A) shows some other important usage activities.
+TABLE II: Comparison of like history across platforms.
+Information Type TikTok Instagram YouTube Minimum expected fields
+User’s usage Watch Y N Y Content Id, ts
+Search Y Y Y Query term, ts
+Comment $N^{*}$ Y Y Comment text, Content Id, ts
+Like Y Y $Y^{g}$ Content Id, ts
+Messages Y Y NA Content, User Id, ts
+Save/Favourite Y Y Y Content Id, ts
+Share(In-app) Y Y NA Content Id, ts
+Share(Across-app) Y N N Content Id, ts
+Interests/Topics Y Y N List of topics
+Time spent $^{*} NN^{*}$ $N^{*}$ Duration/frequency
+User’s content Media Y Y Y Media file/URL
+Text details Y Y Y title
+Location Y Y Y Some place identifiers
+Date time Y Y Y ts
+Device - Y $Y^{g}$ Device model, OS
+Other user interactions $^{*} NN^{*}$ $N^{*}$ Likes, Comments
+Personal details Account details Y Y $Y^{g}$ Username, DOB, Email, Profile photo
+Connections Y Y Y Username, ts
+Login history Y Y $Y^{g}$ IP, ts
+Current devices Y Y $Y^{g}$ User agent
+Current camera NA Y NA Version/type
+Location Y Y $Y^{g}$ Place identifiers
+Account changes - Y N Type, Old, New values, ts
+Advertisements Ads viewed N N Y Content Id, ts
+Personalization $N^{*}$} N N Reasons why the ad was shown
+Access to your data N Y N Which and how (in store visit etc.)
+Miscellaneous Off-platform Y Y N Platform, ts, activity
+Link history - Y - Link, ts
+Cookies - Y - -
+TABLE III: Data Transparency: Overview of shared information (as of Dec, 2024).
+$N^{*}$} denotes the details are found in the app but not in the GDPR dump.
+$Y^{g}$} denotes details are found in google’s DDP, but not in YouTube’s DDP.
+(ts: timestamp).
+3.2 User engagement
+One of the most important categories of information in the DDPs is how a user engages with social media platforms.
+It may include the content that a user watches, searches, or engages with implicitly or explicitly.
+Watch history: Watch history is an ordered list of all the content that a user has watched before requesting the data.
+DDPs of all three platforms contain watch history data of the data subject.
+However, as Table I shows the data provided across the three platforms are widely varying.
+Among these, two points are strikingly different – (a) while YouTube provides watch history for the entire lifetime of a user on the platform, Instagram provides data for at most the last two weeks and TikTok provides data for around six months before the data request.
+These widely varying retention periods for watch history are provided without clear explanations for the criteria used, raising questions about full adherence to Article 15(1)(d) which requires information on data storage periods or the criteria for determining them.
+(b) While TikTok and YouTube provide the watch history all at once, Instagram segregates it into three files based on the type of content, i.e., ads, posts, and videos watched.
+Like history: Similar to watch history, like history is an ordered list of all the contents that a user has liked before requesting the data.
+Table II demonstrates that there is a clear lack of consistency in how platforms currently implement Article 15(3) GDPR.
+The most surprising observation is YouTube does not even include like history in its DDP, despite ‘likes’ clearly constituting personal data processed by the platform and thus potentially falling under the scope of data subjects’ Right of Access.
+We would like to note that although Like history is not there in YouTube DDP, it is included in the Google DDP.
+Time spent on platform: Despite showing such information on their apps, none of the platforms share the time users spend on platforms on a daily basis in the DDPs.
+Only TikTok has recently added a new field in the data, namely ‘Activity Summary’, which enlists the number of videos a user has commented on, shared, or watched until the end.
+Other usage activities on platform:  Table III shows other usage activities.
 Most of the activities, e.g., comment, search, save, share, and writing a message are recorded by all platforms where the features are applicable.
 However, sharing across applications, i.e., when a user shares the video on other social media platforms or copies the link for posting elsewhere, is only shared in TikTok DDPs.
-Also, while both TikTok and Instagram maintain a list of inferred interests of the data subject, YouTube DDPs do not have them.
-3.3 Information regarding user’s content
+Also, while both TikTok and Instagram maintain a list of inferred interests,YouTube DDPs do not have them.
+3.3 User’s content
 While the usage data is about how users consume or behave on video streaming platforms, information regarding their content refers to the content that a user creates and uploads to the platform for others to consume.
@@
 textual captions,
-date, and time information
+date and time information
 TikTok,
@@
 YouTube
-Table 3: Details of user created content shared by platforms.
+TABLE IV: User created content shared by platforms.
 While all three platforms provide a copy of the media and text details, Instagram shares additional information including software used to upload the content, device ID, metadata about the camera, etc.
-(Table 3).
-3.4 Information regarding personal details
-While the two facets discussed above focus more on the content that users consume, engage with, or create, the information regarding personal details is more privacy-sensitive.
-These details contain basic personally identifiable information (PII), e.g., account details (which may include one’s name, phone number, e-mail address, date of birth, profile picture, etc.).
-However, apart from these basic personal details, most of the DDPs also include information about the different devices through which people visit the platforms.
-TikTok and Instagram directly store these details as part of the login information, whereas YouTube does not offer this data in its DDP.
-For YouTube, this data can be accessed only if someone requests their entire Google DDP.
-[^2] One justification may be that one can not sign in to their YouTube account unless they sign in to their Google account on the same device.
-Login history: In light of the prior descriptions, login history contains the list of login activities made by the data subject.
+(Table IV).
+3.4 Personal details
+Distinct from data on content engagement, information regarding personal details is inherently more sensitive.
+These details contain basic personally identifiable information (PII), e.g., account details, one’s name, phone number, e-mail address, date of birth, profile picture, etc.
+However, apart from these basic details, most of the DDPs also include a lot of other sensitive information about the users.
+Login history: Login history contains the list of login activities made by the data subject.
+While TikTok and Instagram provide these details in their DDPs, YouTube does not provide them in its DDP.
+For YouTube, these details are usually found in the Google DDP.
 The login history details are significantly different in TikTok and Instagram DDPs.
-TikTok provides the timestamp, IP, device model, operating system, network type, and carrier provider in its DDP.
-On the other hand, Instagram’s login history data is very detailed.
-Along with providing the above details, Instagram provides (and therefore collects) cookie information, language code, Instagram app version, display properties of the device, hardware identifier, and some internal identifiers.
-Such differences in data collection and sharing practices point to two important possibilities: (a) TikTok might be collecting some of these details and not sharing them with its end-users and, therefore, is not being transparent; (b) Instagram might be collecting more information than necessary, thus becoming more intrusive to privacy.
-Diving deep into these possibilities is beyond the scope of this study, and we would like to explore these directions in future work.
-3.5 Information regarding advertisements
-Advertisement on social media platforms is a prevalent phenomenon, as users are constantly served with ads when they scroll through Instagram, TikTok, or YouTube.
+While TikTok provides the timestamp, IP, device model, operating system, network type, and carrier provider in its DDP, Instagram’s login history data is significantly more detailed.
+Along with the above details, Instagram provides (and therefore collects) cookie information, language code, Instagram app version, display properties of the device, hardware identifier, and some internal identifiers.
+Such differences in the amount of PII shared point to two important possibilities: (a) TikTok might be collecting some of these details and not sharing them with its end-users and, therefore, is not being transparent; or (b) While such detail might be argued as relevant for security purposes, the extent of Instagram’s collection, particularly regarding persistent identifiers and granular device specifics, warrants scrutiny under data minimization principles (Article 5(1)(c)).
+Investigating these is beyond the scope of this study, and we would like to explore these directions in future work.
+3.5 Advertisements
+Users are constantly served with ads when they scroll through TikTok, Instagram, and YouTube.
 However, based on our observations, there is no standardized way of including ads that users have seen in their requested DDP.
-Ads viewed: As mentioned in Section 3.2, Instagram maintains a separate list of ads viewed by users.
-However, the data provided by Instagram only mentions the author/advertiser name and timestamp.
-TikTok, on the other hand, does not demarcate any advertised content that a user has seen in their timeline.
-In contrast, YouTube distinguishes all the ads that a user has encountered from the organic content.
-Also, YouTube provides a link to the ad video along with its title and when it was viewed.
+Ads viewed: While YouTube and Instagram demarcate ads viewed by users, TikTok does not distinguish them in its DDP.
+Among the two platforms that provide ads data, Instagram maintains it in a separate list, whereas YouTube labels ads in its watch history file.
+YouTube provides ads’ title, URL, advertiser details, and timestamp; in contrast, Instagram only shares the advertiser name and timestamp.
 However, neither of the three platforms provides information on ad targeting parameters.
-From the ads viewed details, a user can understand these are the ads they got to watch on their timeline, but why those particular ads were served to them remains an enigma.
-While this information does appear when an advertisement or paid content appears on the three platforms, it is not included in the shared DDPs.
-Ad targeting data sources: The Instagram DDP contains a list of advertisers who have used data about one’s online activity or profile for targeted advertising purposes.
+From the ads viewed details, a user can understand these are the ads they viewed on their timeline, but why those particular ads were served to them remains an enigma.
+This lack of transparency regarding ad targeting could be seen as not fully aligning with the spirit of Article 15(1)(h) concerning meaningful information about automated decision-making, including profiling.
+Ad targeting data sources: The Instagram DDP contains a list of advertisers who have used data about one’s online activity or profile for targeted advertising purposes, specifically to reach them as an identified individual (as opposed to targeting based on general demographics or interests).
 This list also conveys information about how these advertisers obtained users’ data.
-In specific, as per this information, there are three ways in which the advertiser might target a person: (a) custom audience – where the advertiser targets the data subject through a custom audience with a list of customer data (e.g., email addresses, phone numbers, or other identifiers) which the advertiser might have obtained from other third party sources; (b) remarketing custom audience – where the advertiser targets the data subject because (s)he might have interacted with some of the advertiser’s content, visited their website, etc.; (c)  in-person store visit – where advertiser can use location data or check in data because the data subject might have visited one of their physical stores.
-The DDPs of YouTube and TikTok do not provide any such information on advertisers and how they gather data about a particular user.
-Off-platform activities: Another important source of data through which these platforms collect enormous amounts of digital behavioral data is off-platform activities.
+In specific, as per this information, there are three ways in which the advertiser targets a user: (a) custom audience – where the advertiser targets the user through a custom audience with a list of customer data (e.g., email addresses, phone numbers, or other identifiers) which the advertiser might have obtained from other third party sources; (b) remarketing custom audience – where the advertiser targets the user because (s)he might have interacted with some of the advertiser’s content, visited their website, etc.; (c)  in-person store visit – where advertiser can use location data or check in data because the data subject might have visited one of their physical stores.
+YouTube and TikTok DDPs do not provide any such information on advertisers and how they gather data about a particular user.
+Given that ad targeting often involves processing data from various sources, this omission may be a potential violation of Article 15(1)(g) which mandates disclosing such sources and the gathered data.
+Off-platform activities: Another important source of data through which these social media platforms collect enormous amounts of digital behavioral data is off-platform activities.
 This is a list of all the activities in external platforms (outside the source platform) that these platforms have tracked or linked to a user.
-Note that this might be an artifact of the said external platforms’ installed Meta/Google/TikTok pixels.
-A user might have used a Meta or Google login to sign up, other tools embedded in the software development kit of the external platform, or any data-sharing agreement between the external company and any of the platforms we investigate.
-For instance, if a user visits the website or application that also installs one of the tracking techniques of any of the short-format platforms, then the said website can share the user’s activities, such as page view, purchase, search, logins, etc., along with other activities, such as adding a product to the cart, view contents, install app, launch app, etc.
+For instance, if a user visits a website or application that also installs one of the tracking techniques of any of the short-format platforms, then the said website can share the user’s activities, such as page view, purchase, search, logins, etc., along with other activities, such as adding a product to the cart, view contents, install app, launch app, etc.
 Instagram and TikTok DDPs contain off-platform activities of users, whereas YouTube does not have such data.
-Both Instagram and TikTok share the external platform name, event or activity type, and the timestamp in their DDPs.
-The DDPs from the three platforms also provide a wide variety of other details.
-For example, they may contain information on link history, survey, application setting, shopping, etc.
-For more details on this, please refer to the summary of all the discussions above, listed in Table 5 (see Appendix A).
-All three platforms miss one of the important goals set by GDPR Article 15(1)(a), which is to include the purposes of the data processing and collection.
-None of the DDPs include a purpose statement explaining why certain data is collected, nor do they provide a README or manual to help users understand the significance of the data being collected.
-Important takeaways
-[…]
-  The three platforms implement Article 15 very differently.
-The Instagram and TikTok DDPs contain a significantly larger number of data types than the YouTube DDP.
-Instagram shares (and thus collects) more data.
+Both Instagram and TikTok share the external platform name, activity type, and the timestamp.
+The DDPs from the three platforms also provide a variety of other details.
+For example, they may contain information on link history, survey, application settings, shopping, etc.
+For more details, please refer to Table III.
+3.6 Key omissions in current implementation
+While the previous sections bring out some platform specific deficiencies in the content for different categories of data, the GDPR Right of Access requires the implementation to communicate much more than just the data.
+For instance, Article 15(1) mandates the purpose of data collection, the third-party or categories of third-parties to whom such data has been disclosed, the period for which data will be retained, etc.
+Worryingly, all three platforms fail to report this information in accordance with the Right of Access.
+While we acknowledge that some of the information might be available in different policy-related pages of the platform, we posit that it is the obligation of the platform to make this information explicit and accessible.
+Furthermore, none of the DDPs provide any README or manual to help users understand the significance of the different categories of collected data.
+As explained in this section, most often, platforms provide the raw data without providing any metadata or context.
+Without proper context, digesting and making sense of such complex information may be further difficult for common users.
+Important takeaways:
+Despite offering similar services, there are significant inconsistencies in data shared (and thus gathered) by the three platforms.
+Instagram collects disproportionately higher granularity of personal data than TikTok and YouTube.
+There are inconsistencies in the reporting of data within platforms.
+For example, while some details (e.g., URL to the content) are available for one category of data (like history) they are missing in another category (watch history) for Instagram DDP.
+All the studied platforms fail to report purpose, recipients, and retention period of collected data. These shortcomings defeat the purpose of such data protection rights.
+4 Auditing reliability of DDPs
+While the platforms differ in their data collection and sharing practices, as long as they reliably share the information with users it still is acceptable.
+Therefore, in this section, we audit the reliability of DDPs provided by the platforms along the following dimensions.
+Completeness: Evaluates whether the platforms provide a comprehensive record of all user activities or the data is subject to sampling or omissions.
+Correctness: Determines whether the sequence of recorded activities in the DDP accurately reflects the chronological order of events.
+Consistency: Assesses the uniformity of data within same user account upon different requests and across different user accounts.
+4.1 Data collection
+To evaluate completeness and correctness of the data within DDPs, we need access to the ground truth i.e., what are the actual activities done on the corresponding user account.
+Hence, we use sock-puppet accounts to systematically browse and log the behavior of the sock-puppet to generate the ground truth.
+Next, we elaborate on the detailed procedure of data collection through sock-puppet accounts.
+Sock-puppet accounts: We automated video browsing and liking activities using pyautogui [2] to control mouse and keyboard and opencv-python [1] to detect the like button on the desktop screen.
+During the activities, we captured interactions with the browser by downloading the HAR (HTTP Archive) files (HAR is a JSON-formatted file that logs all network traffic between the browser and the server).
+We viewed between 20 to 25 short-format videos in each video browsing session on each of the three platforms.
+Our bot accounts watched each video for a random duration, ranging from 15 to 60 seconds, and they randomly liked some of the videos.
+This approach allowed us to gather accurate ground truth data regarding user activities.
+With this ground-truth dataset, to assess the reliability of the platform provided DDPs, we, then, exercise the GDPR Right of Access from these accounts to get the DDPs from the platforms.
+This process was conducted for nearly thirty video browsing sessions between October 2024 and December 2024 for each platform to avoid any form of stochasticity in observations.
+While the above data enables us to audit the correctness and completeness of DDPs, it does not fully capture the different nuances of consistency.
+For example, multiple time requesting for DDPs in the above approach allows us to audit consistency within an account.
+However, we still can not claim anything about the generalizability of the observation across accounts.
+Therefore, to better assess the consistency of the DDPs, we decided to collect data from real users.
+Real-world users: To collect data from real users, we designed our own data donation website.
+To ensure user privacy, all collected data was anonymized and any PII was discarded on the front-end of our data donation website.
+During the data collection process, we intentionally mandated the donation of specific data categories for each platform.
+For TikTok and YouTube, video browsing history was mandatory.
+In the case of Instagram, the mandatory categories included (a) like history and (b) ads and topics, which consist of posts, ads, videos viewed, and ads clicked.
+Other data categories, such as search history, comments, etc., were optional and users provided them at their discretion.
+Private messages and other sensitive data were not collected.
+We recruited the users on Prolific [25] and employed the platform’s screening features to select the participants who met our eligibility criteria.
+Specifically, participants were required to have an approval rate of at least 98%, a minimum of 100 prior contributions and have regular engagement with Instagram, TikTok and YouTube.
+We also mandated that the users should have at least 90 days of activity on each of these platforms.
+Users were compensated \$5 for the mandatory categories and \$1 for any additional category.
+We collected the data from the same set of users, i.e., all the users were compensated for their data donations for all the three platforms.
+Geo-locations for our data collection: First, we collected data from 10 participants from Germany, France, Italy and Spain – countries within the European Union that have the highest GDP [31].
+While the primary focus of our study is to assess the implementation of Right of Access under the GDPR (which affects residents within EU), to understand the potential ‘Brussels effect’ happening in enabling Right of Access across the world, we collected data from 10 participants each from the UK, Brazil, the USA, and India [^2] For users in India, only Instagram and YouTube data were collected, as TikTok is banned in the country..
+Table VIII in Appendix B shows the user demographics.
+4.2 Audit methodology
 […]
 […]
-4 RQ2(a): How comprehensible are DDPs?
-Here we investigate whether the current implementations of platforms adhere to the requirements Chapter 3 of the GDPR outlines a number of rights of data subjects (i.e., end users), including the right of access (Article 15) [12].
-Before outlining these rights, the GDPR provides a list of requirements for how information should be communicated to end users.
-Specifically, the first paragraph of Article 12 of GDPR reads, “The controller shall take appropriate measures to provide any information referred to in Articles 13 and 14 and any communication under Article 15 to 22 and 34 relating to processing to the data subject in a concise, transparent, intelligible and easily accessible form, using clear and plain language $\dots$ [12].
-As the GDPR does not provide more succinct definitions of these requirements, we adopt two approaches to interpret these terms.
-First, following trends in empirical contract research to employ user surveys for interpretation of legal terms [5], we conduct a large-scale user survey among participants from different European countries to understand their interpretations of some of the requirements.
-Second, we apply the interpretations of the requirements by the European Data Protection Board (EDPB) [14].
-Finally, we evaluate the current implementations, i.e., the DDPs obtained from Instagram, TikTok, and YouTube, against the interpretations by users and the EDPB.
-Participant recruitment: We recruited participants to conduct our survey from the crowd-sourcing platform Prolific[27].
-On Prolific, we hired 100 participants each from Germany, France, Spain, and Italy (400 in total) who had a high approval rate of at least 98% to take part in our survey.
-Our motivation for a general set of respondents stems from the fact that the GDPR entitles general residents in the European Union(EU) with the right to data access, and requirements are intended to benefit these residents.
+[…]
+[…]
+4.3 Observations
+Completeness: As mentioned in Section 3.2, Instagram and TikTok do not provide the entire watch history of a user, since they register on the platform.
+Hence, these DDPs are incomplete and potentially violate Article 15(1)(d).
+Next, we evaluate completeness of the provided data.
+Instagram offers two video browsing feeds: (1) home feed, and (2) reels feed.
+For Instagram’s home feed, all the activities (watching and liking) present in our recorded logs were also present in the requested DDPs.
+However, we observed that videos watched under the reels feed were not present in the DDP, although like activities associated with them were recorded in the DDP.
+This observation raises serious reliability concerns on the shared DDPs of Instagram.
+On TikTok, all browsing and like activities were fully recorded, with a completeness rate of 100%.
+For YouTube, the recorded entries accounted for 99.5% of the activities observed in the HAR dump.
+Correctness: Table V presents the average Jaccard scores across the video browsing sessions for the three platforms.
+For Instagram, there is a drop in the Jaccard scores of browsing, because of changes in usernames, i.e, account username change or account deletion.
+For TikTok, Jaccard scores achieved 100% accuracy across both browsing and like activities.
+For YouTube, the average Jaccard scores are 100% for date, 99.83% for context, and 99.5% overall.
+The slight drop in the overall score is because of the 0.5% missing entries discussed earlier.
+For TikTok and YouTube, the recorded timestamps showed a minimal difference of $\pm$ 5 seconds compared to the HAR dump values for the same activity.
+However, for Instagram, the difference was as high as up to one minute, leading to a reduction in the accuracy for date and overall.
+Thus, all the DDPs showed higher correctness during our audit.
+Platform (Activity) Date Context Overall
+Instagram (Likes) 100% 100% 100%
+Instagram (Browse) 96% 97% 91%
+TikTok (Browse) 100% 100% 100%
+TikTok (Likes) 100% 100% 100%
+YouTube (Browse) 100% 99.83% 99.5%
+TABLE V: Average Jaccard scores across different video browsing sessions. Concerning correctness, TikTok sits at the top, followed by YouTube and Instagram.
+(a) Instagram
+(b) TikTok
+(c) YouTube
+Figure 2: Ratio of entries retained across multiple snapshots for the same account: Instagram (like history), TikTok (video browsing history and like activities), and YouTube (video browsing history).
+Consistency (intra-user):  Figure 2 presents the ratio of overlapping entries between two DDPs to the total number of entries in the earlier DDP for three key aspects, i.e., date, context, and overall.
+For Instagram, despite using the same accounts, we observed variations in the covered duration of the video browsing history across multiple snapshots, ranging from 6 to 13 days.
+Hence, due to limited availability of browsing history, we focus on analyzing the like history.
+Figure 2(a) shows the results for Instagram.
+We observe 6% of the entries are missing from snapshot 3 to snapshot 4.
+On the other hand,  Figure 2(b) illustrates that TikTok provides consistent data across multiple retrieval attempts.
+Lastly, in  Figure 2(c), we observe that 12% of entries were missing for snapshot pairs (1, 3) for YouTube.
+Among these, nearly 62% entries were advertisements, while the other videos got deleted from the platform.
+To sum up, while we found TikTok to be consistent in our audit, Instagram and YouTube failed our consistency audits.
+Consistency (inter-user): However, these observation could be stochastic for being observed in one sock-puppet account per platform.
+To understand the extent to which the distribution of provided data differs across users, we analyzed the cumulative distribution function (CDF) of the durations of video browsing, search, and like history across the platforms.
+Figure 3 presents a comparative view of activity durations across users.
+For Instagram, we observe distinct clusters of users with browsing history durations concentrated around 6 and 13 days (Figure 3(a)).
+This non-uniform distribution has also been observed within the same user upon multiple requests over time.
+On TikTok (Figure 3(b)), video browsing and search history durations exhibit substantial variation.
+Specifically, some end-users had data spanning approximately 180 days, while others had nearly 455 days of recorded video browsing and search activities.
+To ensure a fair comparison, we specifically restricted our analysis to TikTok users who have been using TikTok for at least 450 days.
+This criterion allows us to focus on the subset of users who have had ample time to accumulate browsing and search history.
+As a result, we analyzed data from 49 TikTok users.
+These findings highlight TikTok’s tendency to provide data inconsistently among its users.
+However, such clusters were not observed for like history on both Instagram and TikTok.
+(See Figure 10 in Appendix LABEL:appendix:reliability).
+The analysis of video browsing and search history on YouTube (Figure 3(c)) does not reveal any discernible patterns, suggesting that the platform may provide DDPs more consistently to its end users.
+This observation holds across both the EU and the other regions.
+Inter-user consistency (across EU countries): Such inter-user consistency analysis shows even more interesting patterns across different countries within the EU.
+Instagram users in EU countries typically receive either one week or two weeks of browsing history (see Figure 4(a)).
+In contrast, TikTok displays considerable inconsistency: users within the same countries (Germany and France) often receive different duration of data, despite being part of the same regulatory region (see Figure 4(c)).
+Inter-user consistency (across regions): We observe the ‘Brussels effect’ in terms of users in various other countries (UK, Brazil, USA and India) it is also possible to request and get a copy of personal data.
+Instagram continues to provide users with either one or two weeks of data across regions (see Figure 4(b)), following the same pattern observed across EU countries.
+On the contrary, as shown in Figure 4(d), we find that TikTok shows more uniform behavior in the USA and Brazil (around 180 days of browsing history data is provided) compared to the erratic behavior within EU.
+(a) Instagram
+(b) TikTok
+(c) YouTube
+Figure 3: Plots illustrating the CDF of data duration provided to users for two activities – browse and search history.
+On Instagram, clusters in browsing history are observed at 6 and 13 days, while on TikTok, clusters appear at 180 and 455 days for both search and browse history.
+(a) Instagram (EU)
+(b) Instagram (Regions)
+(c) TikTok (EU)
+(d) TikTok (Regions)
+Figure 4: Plots showing the percentage of users with different browsing history durations.
+For Instagram, users within the same country or region were either given one week or two weeks of data.
+TikTok, by contrast, exhibits greater variability in the EU whereas in the USA, users were provided with the same duration of data.
+Important takeaways:
+TikTok DDPs are the most reliable. While, Instagram failed the completeness and consistency audits for our sock-puppet accounts, YouTube failed the consistency audit.
+Instagram and YouTube exhibit varying degrees of missing data across snapshots. Moreover, Instagram’s DDP is not complete as it does not contain viewed reels.
+There are disparities in the amount of data within a platform’s DDPs across users within EU.
+Instagram and TikTok have clear variations for video browsing and search histories, while YouTube appears to maintain a more consistent data distribution by providing data for the entire history.
+5 Assessing comprehensibility of DDPs
+Despite the presence of the most reliable form of information, the utility of DDPs will be significantly undermined if the communication of such information is not comprehensible to users.
+Therefore, in this section, we investigate whether the current implementations of platforms adhere to the GDPR’s comprehensibility requirements.
+Methodology overview: Per Article 12 of GDPR, DDPs should be in a concise, transparent, intelligible and easily accessible form, using clear and plain language [10].
+As the GDPR does not provide more succinct definitions of these requirements, following trends in empirical contract research  [4], we conduct a large-scale user survey among participants from different European countries to understand their interpretations of some of the requirements.
+Finally, we evaluate the current implementations, i.e., the DDPs obtained from TikTok, Instagram, and YouTube, against the interpretations by participants.
+Participant recruitment: To conduct our survey, we recruited 100 participants each from Germany, France, Spain and Italy (400 in total) who had a high approval rate ($\geq 98\%$) on Prolific [25].
+Our motivation for a general set of respondents stems from the fact that the GDPR empowers general residents in the European Union (EU) with the Right to data access, and requirements are intended to benefit these residents.
 Though, eventually, our plan is to extend the study to all countries within the EU, we chose to start with the countries that have the highest GDP [31].
 We also selected the standard gender breakdown while setting up the surveys.
-In total, we recruited 228 males (57.0%), 161 females (40.3%), 9 self-reported as “other” (2.3%), and 2 participants who preferred not to disclose their gender (0.5%) (See Table 6 in Appendix C for details).
+In total, we recruited 228 males (57.0%), 161 females (40.3%), 9 self-reported as “other” (2.3%), and 2 participants who preferred not to disclose their gender (0.5%) (See Table VIII in Appendix B for details).
 Prior to entering our survey, we presented an online consent form for expressing explicit consent from the participants.
 Upon completion of the survey, we compensated each participant at a rate of £9 per hour, which is recommended by Prolific to be a good and ethical rate of remuneration [24].
-Our survey comprised three components: (a) Awareness of participants about the rights, (b) Interpretation of the requirements by the participants and (c) adherence evaluation.
-Next, we elaborate abut the survey setup and important observations for each of the components.
+Our survey comprised of three components: (a) awareness of participants about the rights, (b) their interpretation of the requirements and (c) adherence evaluation.
+We elaborate on the survey setup and important observations for the three components.
 5.1 Awareness about the GDPR Right of Access
@@
 5.2 Interpretation of the requirements
-Survey setup: To understand the interpretation of requirements by common end users, in the second part of our survey, we showed the participants the paragraph of Article 12 (see above), which mentions the different requirements of DDPs.
+Survey setup: To understand the interpretation of requirements by common end users, we showed the participants the relevant paragraph of Article 12, which mentions the different requirements of DDPs.
 We asked our participants about their interpretation or expectation of the shared personal data in the DDPs to follow any of these requirements.
-Next, we elaborate on the different interpretations of these requirements from the responses of our participants while contextualizing them further with the guidelines of the EDPB [14], which represents the data protection authorities of all EU member states.
-Conciseness Interpretations by the survey participants: We perform thematic analysis of the responses provided by the 400 participants using BERTopic [16].
-The important themes that resonated across the responses spanning all countries are – prioritize important & relevant data, summarize important observations, and present data in a structured way that is easier to navigate.
-However, a fourth theme emerged in the interpretations shared by participants from Italy, France, and Spain, but not by those from Germany.
-Specifically, these participants touched upon the concept of data minimization.
-To this end, some participants mentioned the documents should be shorter, while others mentioned avoiding an overwhelming amount of text.
-In short, while participants from Germany only prioritize the informativeness of the content, other participants prioritize the length of the content as well.
-Interpretations by the EDPB: In this context, the European Data Protection Board interprets concise presentation to be done efficiently and succinctly to avoid information fatigue [14].
-While the above interpretation falls within the interpretations mentioned by our survey participants, the EDPB also asserts that the presentation should differentiate privacy-sensitive information from non-privacy-sensitive information.
-Transparency Interpretation by the survey participants: Unlike the case of conciseness, there are no major differences in how participants from all countries interpret transparency.
+To understand their interpretations of different requirements, we utilized BERTopic [14].
+BERTopic clustered similar interpretations together into ‘topics’; we then manually derived themes based on keywords and representative sentences to understand the underlying abstract nuances.
+Next, we elaborate on the observations in the context of each comprehensibility requirements.
+Conciseness: The important themes that resonated across the responses spanning all countries are – prioritize important & relevant data, summarize important observations, and present data in a structured way that is easier to navigate.
+We observe an interesting difference in the interpretations across countries.
+Participants from Italy, France and Spain mentioned the documents should be shorter, while others mentioned avoiding an overwhelming amount of text.
+In contrast, participants from Germany only prioritize the informativeness of the content irrespective of its length.
+Transparency: Unlike the case of conciseness, there are no major differences in how participants from all countries interpret transparency.
 The interpretation of transparency can be divided into two major aspects.
@@
 Second, and most importantly, some participants also expect the DDPs to mention how the data is processed and for what purpose the data is collected.
-To this extent, the response from one of the participants captures all the notions beautifully: “It (Transparency) means providing clear honest and complete information about how the data is being used.
-There should be no hidden details or surprises for the individual.
-People should easily understand what data is being collected why its being collected who will access it and how long it will be kept.” At this point, we would also like to note that as per the participants’ interpretation, conciseness and transparency are two competing requirements, i.e., to satisfy one, the other possibly has to be violated.
-Interpretation by the EDPB: While the EDPB draws most of its interpretations from Recital 39 of the GDPR, it interprets transparency as allowing data subjects to determine the scope and consequences of the concerned data processing [14].
-To this end, the EDPB calls for data controllers (i.e., platforms) to separately spell out the most important consequences and risks that the data processing may entail.
-Intelligibility Interpretation by the survey participants: Among all the requirements provided by the GDPR, the most subjective one is that of intelligibility.
-However, almost all the participants explicitly interpret it to be easy to understand and link it closely with the language being straightforward and without any technical and/or legal jargon.
+Conciseness vs.
+Transparency: As per participants’ interpretations, conciseness and transparency requirements are at odds with each other.
+Therefore, any DDP representation may not satisfy these requirements simultaneously, i.e., to satisfy one, the other has to be (at least partly) disregarded.
+Intelligibility: All the participants explicitly interpret it to be easy to understand and link it closely with the language being straightforward and without any technical or legal jargon.
 Participants from all countries also mentioned that intelligible data should have a clear explanation of how and why a certain piece of information is collected.
-The participants also noted some interesting expectations in this regard.
-For instance, participants from Germany mentioned that platforms shall share their DDPs in one’s language of comfort (native language or other language of choice).
-Moreover, participants from Italy stated that platforms should share their DDPs in the form of elegant visualization.
-Interpretation by the EDPB: The EDPB describes intelligibility to be a representation that should be understood by an average member of the intended audience.
-In addition, they also mention that it is closely linked to the requirement to use clear and plain language.
-Clear and plain language Interpretation by the survey participants: Understandably, most of the participants interpret plain and simple language to be words that are easily understandable and unambiguous.
+To this end, participants also noted some interesting expectations.
+For instance, participants from Germany mentioned that platforms should share DDPs in one’s language of choice.
+Moreover, participants from Italy stated that platforms should share their DDPs in the form of an elegant visualization.
+Clear and plain language: Most of the participants interpret plain and simple language to be words that are easily understandable and unambiguous.
 Some participants elaborate on the nuance by stating that legal and technical terms should be avoided or at least explained in a simpler way.
-To this end, one clear observation from our survey responses is that participants do not distinguish between intelligibility and language.
-Interpretation by the EDPB: Along the lines of the interpretations mentioned above, the EDPB interprets this requirement as being definitive and should not be phrased in ambivalent terms or leave room for different interpretations [14].
-Accessibility Interpretation by the survey participants: Most of our participants interpret accessibility as being easily available, i.e., it is easy to find the place to request the data and to access the data.
-However, there are certain interesting and valid interpretations which are worth mentioning.
-Some participants mentioned that the data needs to be available in a format so that it can be opened without any sophisticated software or tools.
-At the same time, the timeliness at which the data is provided upon request is another interesting aspect touched upon by some of our participants.
-Another interesting interpretation was the data should be available to everyone, including people with visual impairments by using accessible design features like screen readers or alternative text making sure the users can access data without any barriers.
-Interpretation by the EDPB: The EDPB’s interpretation of accessibility also follows from the above popular responses from the survey participants, wherein they say that it should be immediately apparent to end-users as to where and how the DDP can be accessed [14].
-Our observations on accessibility: Despite agreeing with all of the above interpretations, we observe a few differences across the platforms because accessibility is the most objective property for evaluation.
-In terms of ease of request, from the content page of a user’s account, one needs to click 6, 6, and 10 times on TikTok, YouTube, and Instagram, respectively, to be able to request the data in the human-readable format.
-Notice that, while on Instagram, a user needs to click more, the platform provides many options to the end-user, such as the time duration that their DDP should contain.
-Moreover, based on our anecdotal observations, the turnaround time for TikTok upon request was instantaneous during this research study.
-For Instagram and YouTube, it often takes 10 to 15 minutes for the DDP to become available.
-However, all the turnaround times are within the GDPR-prescribed maximum duration of one month.
-Further, one can download one’s data on TikTok and Instagram from both their applications and web versions, but YouTube shares the DDP with its end users by sending an e-mail with a URL to the DDP.
-All these platforms provide DDPs in a compressed format, which makes it harder for the end-users to open them on a smartphone.
-At the same time, an end user needs to have a browser or text reader to open the HTML (YouTube and Instagram) and TXT (TikTok) DDPs.
-In summary, while the data request and download are relatively accessible in the current setting, reading the data is still complicated for the average end user.
-At this point, we would also like to note that we are not aware of any special provisions that platforms may have for people with visual impairments to access the data.
+To this end, we observe that participants do not distinguish between intelligibility and language requirements.
+Accessibility: Most of our participants interpret accessibility as being easily available, i.e., it is easy to find the place to request the data and to access the data.
+However, there are certain interesting and valid interpretations that are worth mentioning.
+Some participants mentioned that the data needs to be available in a format so that it can be opened without any sophisticated software.
+Also, the timeliness with which the data is provided upon request is an interesting aspect mentioned by participants.
+Another important interpretation was to make sure that users with visual impairments should be able to access their data using features like screen readers.
+Interpretations by EDPB: To understand whether participants’ interpretations align with that of legal practitioners, we refer to the European Data Protection Board’s (EDPB) [12] guidelines which is to be followed by platforms while implementing the requirements.
+Along with agreeing on most of the above interpretations, EDPB also adds some more nuanced interpretations in its guidelines.
+While both participants and EDPB interpret concise presentation to be succinct presentation to avoid information fatigue, EDPB also asserts platforms should differentiate between privacy sensitive and non-sensitive information in DDPs.
+Similarly, in addition to the mentioned interpretation of transparency, EDPB also asks platforms to clearly mention the most important consequences and risks that each data processing category may entail.
+Much like the tension we found between participants’ interpretation, even EDPB recommendations reflect such tensions between conciseness and transparency.
+For intelligibility, EDPB completely agrees with participants’ interpretations by linking it to the usage of clear and plain language which could be understood by an average member of the intended audience.
+Along with understandability, for the interpretation of clear and plain language, EDPB recommends using definitive language that does not leave any room for different interpretations.
+Finally, for accessibility, EDPB’s recommendations completely match that of our surveyd participants.
 5.3 Adherence evaluation
-Survey setup: In the final part of the survey, we presented our participants with the information shared in the DDPs of Instagram, TikTok, and YouTube in an anonymized and randomized format.
-We showed different categories of content and the whole DDP to our participants.
-The different categories of content that we shared include data from all of the categories of information that we elaborated in Section 3.
-Table 7 (refer Section C.1) summarizes the specific thirteen categories that we evaluated.
-[^3] We should point out that YouTube does not provide like history, autofill, location, and off-platform activity.
-Hence, for these categories, we asked our participants to evaluate between Instagram and TikTok.
-Note that, for this comparative study, we used DDPs that are shared in a human-readable format by all three platforms.
-Instagram and Tiktok share the human-readable version in HTML and TXT formats, respectively.
-However, YouTube DDP only contains watch history, search history, personal information, and login history in HTML format.
-All other categories of data are found in CSV or JSON formats.
-Since the primary goal of this study is to compare the contents, we showed them consistently in a $<category,value>$ pair in our survey.
-We acknowledge that the way data is shown to users may affect some of the results.
-However, in this study, we restrict ourselves to the evaluation of what content is shared rather than how it is shown.
-We plan to explore the visualization aspects in future studies.
-For each category, we showed participants the content shared by the three platforms side by side and asked them which platform’s data representation they found to be more concise/intelligible/transparent/in clear and plain language.
-[^4] We exclude accessibility from the evaluation because it pertains to getting access to request and download DDPs more than to their representations.
-Also, we stated our observations about this requirement in Section 4.2.
-Figure 8 (see Section C.1) depicts a representative example of one of the questions that we used in our survey.
-The first preference in Figure 8 means the most preferred choice of data representation of a participant, whereas the third preference means their least preferred choice of data representation.
-We posed this question as a multiple-choice grid, so each participant had to provide an ordered preference for the three representations.
-For example, if a participant rates Instagram’s representation as their first preference, then they cannot rate any of the other platforms as their first preference.
-Note that the final part of the survey consisted of 14 (13 categories + overall DDP) $\times$4 evaluations where participants gave us an ordered preference.
-Hence, to avoid participant fatigue, we asked 50% of our participants to evaluate intelligibility and transparency and asked the other half to evaluate conciseness and language properties.
-Hence, 200 participants answered each of the evaluation questions.
-Observations
-Next, we report our findings on the adherence evaluation of four requirements.
-Language: Figure 2 shows the percentage of first preference votes elicited by the representation of each of the platforms for the four requirements that we surveyed.
-In the evaluation of clarity and simplicity of language, Instagram won the first preference votes in eleven out of the thirteen data categories (Figure 2(a)).
-In contrast, for the remaining two categories, participants preferred the language in which YouTube presents its search and watch histories.
-To understand the overall preference of participants across the board, we weight every first preference vote with a score of 3, second preference vote with a score of 2, and third preference vote with a score of 1.
-For categories, where we are comparing only two platforms, we assigned scores 3, 1 for first, second preference respectively.
-Then, we evaluate an average preference score for each of the platforms based on participants’ preference orderings.
-Figure 3 shows the average preference scores of each of the platforms in each category of data.
-Based on the average preference scores, apart from watch, search histories and location, the gap between preference toward the language of Instagram’s data representation is way more than that of TikTok and YouTube (Figure 3(a)).
-Across the different categories of data, the participants’ average preference score for Instagram’s language is 2.48, whereas TikTok and YouTube fail to reach an average preference score of 2.
-Apart from the thirteen categories of data, we also asked our participants to evaluate the overall DDP from the three platforms.
-When participants were asked to evaluate the overall DDP, 55% of them preferred the language of Instagram’s DDP over the other two.
-TikTok and YouTube DDPs are preferred by 25% and 20% of the participants, respectively.
-Intelligibility: Similar to Language, Instagram won the first preference vote for eleven out of the thirteen categories evaluated (Figure 2(b)).
-Among these, except for devices and location categories, nearly more than 75% of the participants preferred Instagram’s data representation to be the most intelligible.
-The average preference scores almost follow the same trend as that of language, where the Instagram DDP’s intelligibility (2.53) is substantially higher than that of YouTube (1.75) and TikTok’s (1.64) (Figure 3(b)).
-For the evaluation of the overall DDP, 63% of the participants identified the Instagram DDP to be the most intelligible.
-TikTok and YouTube were chosen as the first preference by a considerably smaller proportion of our participants, nearly 18% each.
-Transparency: YouTube won the first preference for five categories, where over 70% of the participants chose it for watch history, search history, and login, followed by devices and content with 55% and 46%, respectively (Figure 2(c)).
-For the remaining eight out of the thirteen categories, Instagram won the first preference of the participants in a large percentages.
-We find that the average preference scores of Instagram and YouTube are comparable (Instagram: 2.23, YouTube: 2.15), whereas TikTok (1.66) has lower score (Figure 3(c)).
-In the evaluation of the overall DDP, Instagram is preferred more compared to the other two platforms, with nearly 46% of the participants choosing it to be the most transparent.
-On the other hand, YouTube is preferred by 35% of the participants, and the remaining 19% selected TikTok.
-Conciseness: Based on Figure 2(d), we found that in nine out of thirteen categories of data that we surveyed, respondents preferred TikTok’s data representation as the most concise, whereas Instagram’s data representation is found to be the most preferred for the remaining four categories.
-While TikTok’s representation is preferred substantially in the location, autofill, and off-platform categories ($\geq$60first preference votes), Instagram’s representation is preferred more for likes history and contents ($\geq$65first preference votes).
-Based on the average preference scores for conciseness, we see that the representations are found to be very comparable in most of the data categories (Figure 3(d)).
-The average preference scores of Instagram and TikTok are very comparable (Instagram: 2.12, TikTok: 2.15), whereas that of YouTube is considerably lower (1.60).
-While evaluating the overall DDP, 43% of the participants found Instagram’s DDP to be more concise, followed by Tiktok and YouTube with 35% and 22%, respectively.
-Variations across countries: We do not observe qualitative variation across the four countries for three requirements – language, intelligibility, and transparency.
-However, for conciseness, participants from Germany predominantly preferred Instagram, with over 50% of the participants selecting it as their first choice across most categories.
-In contrast, participants from other countries primarily selected TikTok as having the most concise representation for all the categories, except likes, connections, and content.
-To understand this, we manually examined how participants across different demographics interpret conciseness.
-We found that, except for participants from Germany, most demographics have an interpretation of conciseness as: (1) minimizing document length, (2) avoiding information overload etc.
-This difference in interpretation is reflected in how our participants evaluated the data representations: while participants from Germany gave more priority to content, other participants gave more priority to data categories having less content.
-(a) Clear and Plain Language
-(b) Intelligible
-(c) Transparency
-(d) Conciseness
-Figure 2: Percentage of first preferences across all categories and for the entire DDP.
-Instagram is the top choice for most categories across all requirements, except for conciseness.
-For watch and search histories, YouTube is the most preferred platform in transparency, intelligibility, and language.
-(a) Clear and Plain Language
-(b) Intelligible
-(c) Transparency
-(d) Conciseness
-Figure 3: Comparison of 13 different data categories, including an evaluation of overall DDPs for the four requirements. Instagram achieves the highest overall DDP score for all the four requirements.
-Important takeaways
-[…]
-  We found participants’ interpretation of the requirements are in line with that of European Data Protection Board.
-However, there seems to be conflicting interpretations of conciseness and transparency among both stakeholders.
-[…]
-5 RQ2(b): Improving DDPs comprehensibility
-Figure 4: Comparison of watch history representation in our proposal and that of YouTube’s current DDP implementation.
-(a) Clear and Plain Language
-(b) Intelligible
-(c) Transparency
-(d) Conciseness
-Figure 5: Percentage of votes for the proposal and the earlier winning platform across different categories for each requirement.
-Our proposal emerged as the top choice in all fields for the language and intelligible requirements.
-In Section 4, we observed that none of the platforms adhere to the prescribed requirements of the GDPR properly across all data categories, highlighting the need for better representations.
-Motivated by this, here we utilize the different interpretations we elicit from survey participants to come up with novel recommendations that lead to better representations.
-5.1 Methodology
-To come up with an actionable recommendation for data representation, we take the following steps.
-Step 1: Summarize the interpretations: As a first step, for each of the requirements, we collect the interpretations from our surveyed participants and that of the EDPB [14].
-Next, we feed this entire set of 401 interpretations to a large language model (LLM) – Gemini 2.0 Flash [10]– using the Google AI Studio platform.
-We ask the LLM to generate a set of recommendations that encapsulate the most prevalent interpretations by summarizing the above responses.
-For reproducibility purposes, we make the prompts we used and the responses in  Appendix B.
-Note that in this approach, we include the interpretations of not only the data subjects, but also of data protection authorities (through the interpretation of EDPB).
-Step 2: Generate recommended data field for each data category: In the second step, by providing the top recommendations for each of the requirement, we ask the LLM to make some assumptions.
-Specifically, we state that we are data engineers working in a short-format video platform operating in the European Union.
-Moreover, an end user requests us to provide them with their personal data for a given data category.
-Then, we ask the LLM to let us know of the specific data fields that we should provide to the end user.
-Subsequently, the LLM provides us with a set of recommendations for data fields of information in the queried category in the form of a JSON object (available in Appendix B).
-Step 3: Manual inspection by researchers: In the final stage, three researchers who are part of this study and are conversant with DDPs of a whole array of platforms assess these recommendations.
-The primary goal of their assessment is to remove any instance of hallucination in the form of potentially irrelevant data fields that may have been recommended by the LLM.
-Next, given their expertise in understanding data donations in other platforms like Netflix, Prime video etc., we ask them to suggest any additions or removals to the recommended fields.
-By taking these steps, we ultimately come up with the final proposal of recommendations.
-Note that while Step 1 introduces the end users’ and data protection authorities’ interpretations, this step also instils the common interpretations of tech experts on other platforms and researchers who want to unravel the dynamics of platforms using these DDPs.
-An example of the recommended representation for watch history and that of the most adhering representation among the existing one’s (YouTube’s) is shown in Figure 4.
-Note that apart from the details about the video watched, the proposal also shares the duration of the video and what fraction of it was watched by the user and on which device and network.
-5.2 Evaluation of the proposed representation
-To evaluate our proposed representation, we conducted another survey among 200 participants (who have not participated in the earlier survey mentioned in Section 4) from Germany, France, Spain, and Italy.
-We follow the identical participant recruitment strategy as elaborated in Section 4.
-Table 8 (See Appendix D) shows the user demographics.
-During the survey, we showed the participants two representations for the evaluated categories: (a) proposed representation: which we generate from the steps mentioned in Section 5.1, (b) winner representation: for each data category, the representation among Instagram/TikTok/YouTube which won the first preference along more requirements (i.e., the representation which participants suggested to be more adhering to the GDPR).
-We anonymize and randomly present each pair of representations to the survey participants.
-Then, we ask the participants to select one of the two representations that they find to be more intelligible/transparent/concise/in clear and plain language.
-Observations We compare the representations of the data category fields of our proposal and that of the best among the existing short-format video platforms for 13 different categories across four GDPR requirements.
-[^5] Notice that the data categories are slightly different than the evaluations done in Section 4.3.
-We are not considering personal information, as there were no suggestions for adding or removing any fields.
-Instead, we include activity summary, which is provided only by TikTok, for comparison.
-Figure 5 shows the percentage of votes for the two representations obtained in the 52 evaluations (13 data categories $\times$ 4 requirements), that we surveyed.
-We observe that the participants preferred our proposed representation in as many as 44 categories.
-However, their preferences vary for different requirements.
-Language and intelligibility: For clarity and simplicity of language and intelligibility of the representation, our proposal gets upwards of 67% of the votes in most of the categories (except language in comments and intelligibility in autofill categories) suggesting our proposed representation is both intelligible and in clear and plain language.
-Transparency: Our proposal is voted to be more transparent than the best of the existing short-format video platforms in 11 out of the 13 categories.
-Apart from search history, in all the other winning 10 categories, our representation gets upwards of 75% of the votes.
-In watch history, the split between our proposal and that of YouTube’s DDP is 47% and 53%, respectively, indicating that both representations are of comparable quality.
-However, in autofill information category, Instagram’s data representation gets nearly 93% of the votes from our survey participants.
-However, we argue that having less autofill information is indicative of less privacy-intrusive data collection.
-While our proposed recommendation includes only the email address, name, address, and phone number, that of Instagram contains much more sensitive information.
-Conciseness: Notice that in Section 4.2, we mentioned conciseness and transparency, and to some extent intelligibility, are mutually competing requirements.
-We observe the same conflict in the results of this survey.
-Unsurprisingly, our proposal was preferred to be the most concise representation in merely 7 out of the 13 categories.
-Even then, in watch history, search history, login history, and device details – our proposed representation was found to be more concise by upwards of 75% of the participants.
-It is also worth mentioning that, much like the survey presented in Section 4.3, the participants from Germany have different preferences compared to those from other countries.
-Except for autofill information, over 66% of German participants preferred our proposed representation in all other categories.
-Despite our proposed recommendation coming out as a significant improvement over any of the existing data representations, achieving this was not the main objective of this section.
-Our primary intention of this section has been to convey how, relatively easily, one can come up with more effective data representations by taking into account the interpretations of different stakeholders.
-This potentially indicates that currently, platforms are probably operating in silos without much deliberation and sincere effort into implementing Article 15 of the GDPR.
-At the same time, because of the mutually conflicting nature of the requirements, it is harder than anticipated to satisfy all the requirements with a single representation.
-Important takeaways
-[…]
-  Taking into account the interpretations of different stakeholders (e.g., end users, legal scholars, technical experts, and researchers) can offer a significant improvement to the data representations across different categories.
-[…]
-[…]
-6 RQ3: How reliable are the DDPs?
-In this section, we investigate how reliable the DDPs are.
-We first define the requirements for quantifying the reliability of DDPs.
-Next, we outline the data collection employed to evaluate their reliability.
-Finally, we conduct a comprehensive assessment of the reliability of DDPs for the three platforms.
-6.1 Requirements for reliability
-86% of researchers who participated in a Data Donation Symposium reported encountering problems with the DDPs provided by various platforms [34].
-The primary concerns identified include: (1) incompleteness, such as missing data types or limited time ranges, and (2) inaccuracy, such as wrong timestamps, incorrect ordering which renders the data unreliable to both end users and researchers.
-Therefore, due to the popularity of using DDPs to conduct research, there is a pressing need to systematically assess how reliable DDPs are.
-Hence we evaluate the reliability of DDPs on three requirements.
-Completeness: Evaluates whether the platforms provide a comprehensive record of all user activities or whether the data is subject to sampling or omissions.
-Correctness: Determines whether the sequence of recorded activities in the DDP accurately reflects the chronological order of events and whether any wrong entries are present.
-Consistency: Assesses the uniformity of data across users and within users across different retrieval instances.
-6.2 Experimental setup
-We collect two types of data: (a) from real-world users, and (b) utilizing bot accounts.
-Real-world users Inspired by prior work, we designed our own data donation website, which is privacy preserving and allowed us to obtain DDPs from real-world users [38].
-To ensure user privacy, all collected data was anonymized and any PII was discarded on the front-end of our data donation website.
-During the data collection process, we intentionally mandated the donation of specific data categories for each platform.
-For TikTok and YouTube, video browsing history was mandatory.
-In the case of Instagram, the mandatory categories included (a) like history and (b) ads and topics, which consist of posts, ads, videos viewed, and ads clicked.
-Other data categories, such as search history, comments, etc., were optional and users provided them at their discretion.
-Private messages and other sensitive data were not collected.
-Users were compensated based on the categories and amount of data they donated.
-The maximum payment offered was \$16 for TikTok, \$9 for YouTube, and \$14 for Instagram data donations.
-Data collection for TikTok and YouTube took place between August and September 2024, while for Instagram, it was conducted between December 2024 and January 2025.
-We collected data from the same set of 33 users across all three platforms.
-17 (51.5%) male and 16 (48.5%) female participants donated their data.
-Bot accounts We automated the process of video browsing and video liking activities using pyautogui [2] to control mouse and keyboard and opencv-python [1] to detect the like button on the desktop screen.
-This approach differs from traditional browser automation [7], as we directly controlled the desktop environment instead of interacting with the browser via scripts.
-We captured interactions with the browser by downloading the HAR (HTTP Archive) files (HAR is a JSON-formatted file that logs all network traffic between the browser and the server).
-We viewed between 20 to 25 short-format videos in each video browsing session on each of the three platforms.
-Our bot accounts watched each video for a random duration, ranging from 15 to 60 seconds, and they randomly liked some of the videos.
-This approach allowed us to gather accurate ground truth data regarding user activities.
-The process was conducted for nearly thirty video browsing sessions between October 2024 and December 2024 for each platform.
-To conduct our analysis, we collected multiple DDP snapshots from each of these bot accounts.
-6.3 Observations
-Completeness: As mentioned in Section 3.2, Instagram and TikTok do not provide the entire watch history of a user, since they register on the platform, without any explanations.
-Hence, these DDPs are inherently incomplete and potentially violate Article 15(1)(d).
-However, with whatever data they share, to evaluate completeness, we utilize the data collected from the bot accounts.
-To this end, we compared the number of activities recorded in the HAR dump with those in the DDP for each video browsing session.
-We used the DDP that was closest in time to the corresponding HAR dump.
-Instagram offers two video browsing feeds: (1) home feed, and (2) reels feed.
-Interestingly, we found that videos watched under the reels feed were not present in the DDP, although like activities associated with them were recorded.
-For Instagram’s home feed, the percentage of recorded entries in the DDP for browsing videos and likes is 100%.
-On TikTok, all browsing and like activities were fully recorded, with a completeness rate of 100%.
-For YouTube, the recorded entries accounted for 99.5% of the activities observed in the HAR dump.
-Correctness: After assessing the completeness of DDPs, we evaluated its correctness using the same data from bot accounts, by checking whether the order of contexts was maintained and whether any arbitrary entries were present in the DDP or not.
-To achieve this, we measured the Jaccard similarity for the following aspects between HAR and DDP for each activity type.
-$\bullet$ Date: Consistency of recorded timestamps. We compared them up to the minute level to allow for slight variations.
-$\bullet$ Context: Consistency of video IDs or author IDs.
-$\bullet$ Overall: A combined evaluation of date and context.
-Table 4 presents the average Jaccard scores across the video browsing sessions for the three platforms.
-For Instagram, there is a drop in the Jaccard scores of browsing, because of changes in usernames, i.e, account username change or account deletion.
-For TikTok, Jaccard scores achieved 100% accuracy across both browsing and like activities.
-For YouTube, the average Jaccard scores are 100% for date, 99.83% for context, and 99.5% overall.
-The slight drop in the overall score is because of the 0.5% missing entries discussed earlier.
-For TikTok and YouTube, the recorded timestamps showed a minimal difference of $\pm$ 5 seconds compared to the HAR dump values for the same activity.
-However, for Instagram, the difference was as high as up to one minute, leading to a reduction in the accuracy for date and overall.
-Platform (Activity) Date Context Overall
-Instagram (Likes) 100% 100% 100%
-Instagram (Browse) 96% 97% 91%
-TikTok (Likes and Browse) 100% 100% 100%
-YouTube (Browse) 100% 99.83% 99.5%
-Table 4: Average Jaccard scores across different video browsing sessions for the three platforms. Concerning correctness, TikTok sits at the top, followed by YouTube and Instagram.
-(a) Instagram
-(b) TikTok
-(c) YouTube
-Figure 6: Ratio of entries retained across multiple snapshots for the same account: Instagram (like history), TikTok (video browsing history and like activities), and YouTube (video browsing history).
-Consistency (across snapshots): Next, we determine whether a particular account receives consistent data when making multiple requests from the same platform.
-Towards this goal, we analyzed differences between multiple DDPs obtained from bot accounts.
-We examined whether the entries present in an earlier snapshot were retained in a later snapshot.
-For Instagram and TikTok, the snapshots are collected with a gap of 1 week.
-For YouTube, we use a gap of one month.
-Figure 6 presents the ratio of overlapping entries between two DDPs to the total number of entries in the earlier DDP for three key aspects, i.e., date, context, and overall.
-For Instagram, despite using the same bot accounts, we observed variations in the covered duration of the video browsing history across multiple snapshots, ranging from 6 to 13 days.
-Hence, due to limited available video browsing history, we solely focus on analyzing the like history.
-Figure 6(a) shows the results for Instagram.
-We identify some missing entries in the later snapshots.
-Specifically, 6% of the entries are found to be missing from snapshot 3 to snapshot 4.
-Figure 6(b) illustrate the results for video browsing history and like activities on TikTok.
-Based on our findings, we observe that TikTok provides consistent data across multiple retrieval attempts.
-Lastly, in  Figure 6(c), we observe that approximately 12% of entries were missing for snapshot pairs (1, 3) for YouTube.
-Among these, nearly 62% entries were advertisements, while the other videos got deleted from the platform itself.
-Consistency (across users): To assess whether all users receive a similar amount of data, we analyzed video browsing history, search history, and like history obtained from real-world users across the three platforms.
-First, we measured the duration of data that was contained in each activity history by calculating the time difference between the earliest and the latest entries in the respective activity history lists.
-For Instagram, we defined browsing history as the combination of ads viewed, posts viewed, and videos watched, while we included both keyword or phrase searches and user searches in the search history.
-To understand the extent to which the distribution of provided data to the users differs, we analyzed the cumulative distribution function (CDF) of the durations of video browsing, search, and like history across the platforms.
-Figure 7 presents a comparative view of activity durations across users.
-For Instagram, we observe distinct clusters of users with browsing history durations concentrated around 6 and 13 days (Figure 7(a)).
-This non-uniform distribution indicates that some Instagram end users receive significantly different amounts of data.
-Similarly, on TikTok (Figure 7(c)), video browsing and search history durations exhibit substantial variation.
-Specifically, some end-users had data spanning approximately 180 days, while others had nearly 450 days of recorded video browsing and search activities.
-To ensure a fair comparison, we specifically restricted our analysis to TikTok users whose earliest recorded activity occurred before March 2023.
-This criterion allows us to focus on the subset of users who have had ample time to accumulate browsing and search history.
-As a result, we analyzed data from 23 TikTok users.
-These findings highlight TikTok’s tendency to provide data inconsistently among its users.
-However, such kind of clusters were not observed for like history on both Instagram and TikTok.
-The analysis of video browsing and search history on YouTube (Figure 7(d)) does not reveal any discernible patterns, suggesting that the platform may provide DDPs more consistently to its end users.
-(a) Instagram - Browse, Search
-(b) Instagram - Likes
-(c) TikTok
-(d) YouTube
-Figure 7: Plots illustrating the CDF of data duration provided to users for three activities – search, browse, and like history.
-On Instagram, clusters in browsing history are observed at 6 and 13 days, while on TikTok, clusters appear at 180 and 455 days for both search and browse history.
-Important takeaways
-[…]
-  In terms of correctness and completeness, TikTok provides the most reliable DDPs. Also, its DDPs are consistent across multiple retrievals.
-[…]
-[…]
-7 Concluding discussion
-Summary of the insights: To the best of our knowledge, ours is the first systematic attempt to evaluate the implementation of the right of access (Article 15) across platforms offering similar services.
-Our detailed evaluation of Instagram, TikTok, and YouTube, focusing on their DDPs, reveals that Instagram and TikTok share more data categories, while YouTube’s shared data is limited.
-The only underlying similarity that cuts across platforms is that of non-disclosing the purpose of data collection and processing.
-While Instagram’s data representation is deemed most comprehensible, TikTok’s is found to be the most reliable.
-Finally, a proposed data representation, leveraging interpretations of different important stakeholders, is found to be preferred by participants across multiple countries along the different requirements.
-Limitations of the current work: Like any other work it has its own share of limitations.
-The survey population could have been more diverse and representative of common users on the street.
-However, note that our work is comparative by nature and we do not expect that the choice of participants will impact the relative rankings and comparisons presented across the paper.
-Further, our study solely focuses on data shared by platforms upon request and not the data that is actually collected from the users.
-Including this information may further enrich the quality of the work.
-Recommendations for stakeholders: Platforms should improve the consistency among different categories of data that they share with end-users.
-The way different related categories of data are being shared differently (e.g., like vs.
-browsing history on Instagram) raises concerns about the lack of coordination within the platform while implementing the regulations.
-The other significant room for improvement for platforms is increasing the comprehensibility of the shared data.
-As we show in our work if we consider the interpretation of different stakeholders, the resulting representations can already be a significant improvement on the current representations.
-The current differential and ineffective implementation not only affects end-users’ rights of access but also affects their rights to port their data (Article 20).
-Such rights of users will only be useful if data protection authorities provide more detailed technical specifications and/or standardization for implementation.
-As Section 5 shows involving different stakeholders may further improve the standard of implementations.
-Moreover, authorities need to enforce these requirements–otherwise platforms may not have any ramifications to worry about if they do not comply with them.
-To this end, the current study also indicates how effective enforcement of the GDPR rights may benefit from computer science research by relying on audit strategies to observe compliance at scale.
-Acknowledgments
-We would like to extend our sincere gratitude to Prof.
-Christoph Engel (Max Planck Institute for research on Collective Goods) and Prof.
-Cristoph Sorge (University of Saarland) for their valuable feedback during the course of this work.
-Additionally, Ingmar Weber is supported by funding from the Alexander von Humboldt Foundation and its founder, the Federal Ministry of Education and Research (Bundesministerium für Bildung und Forschung).
-8 Ethics considerations
-We placed the highest priority on ensuring that our research project adheres to standard ethical guidelines.
-To this end, and before initiating our study, we acquired approval from the Ethical Review Board (ERB) of our University.
-Before surveying a participant or obtaining their DDPs, we provided them with an online consent form and asked them to meticulously review the outlined terms in the consent form and explicitly provide us with their consent to these terms.
-We refrained from collecting any PII from our participants who took part in our surveys and donated us their DDPs.
-Specifically, we removed any available PII inside the DDPs at the front-end and also anonymize them before they are sent to our back-end server.
-Furthermore, according to the stated terms and conditions in the ERB approval, we took stringent measures to protect the privacy of our participants.
-In particular, we would not share their donated anonymized DDPs with any third parties, and we would completely remove these DDPs within a specified time after the conclusion of our study.
-When conducting our surveys, we ensured that the questions did not pose any potential risks to our participants.
-Specifically, when surveying our participants on different representations of the DDPs, we filled out data categories based on sample DDPs that our research team obtained from their own personal accounts on the studied platforms.
-Moreover, we checked the contents of these data categories and verified that they do not endanger our participants in any way.
-We utilized bot accounts that exclusively interacted with publicly available content on the platforms to carry out one part of our experiments in Section 6.
-We took multiple steps to minimize the inadvertent consequences of running these bot accounts.
-To this extent, we intentionally distributed the incurred overhead of our experiment over multiple short browsing sessions to prevent causing any sort of disruptions to the short-format video platforms.
-Specifically, on a single day of our experiment, a bot account had only one video browsing session and viewed up to a maximum of 25 short-format videos.
-Moreover, although our bot accounts tangibly engaged with some videos by liking them, we used only three bot accounts across all the platforms, so we do not foresee any potential harm resulting from these activities.
-Furthermore, although the Terms of Service of the studied platforms prohibit the usage of automated bots, the benefits of our study outweigh any potential risks associated with the activities conducted by our bot accounts.
-Finally, we affirm that we fully conformed to ethical research standards throughout our work [28].
-This includes but is not limited to, presenting our findings in aggregate form and guaranteeing the anonymity of all our participants.
-9 Open science
-We comply with standard open science guidelines to the fullest extent permitted by the terms and conditions outlined in the ERB approval that we obtained.
-To this end, we publicly share multiple materials that we utilized to carry out our research project in References.
-Specifically, we provide details on the surveys which we conducted in Appendix B.
-These details include the exact questions and choices that we gave to our participants.
-Moreover, we present details on the prompts we fed to the LLM and the responses that we got from the LLM also in Appendix B.
-As previously stated in Section 8, according to the terms outlined in our approved ERB proposal, we would remove all the DDPs that we collected in Section 6 from our database after the completion of our research project.
-Therefore, we cannot make these anonymized DDPs publicly available.
-Appendix A Content of DDPs provided by YouTube, TikTok, and Instagram.
-Table 5 presents an overview of the content included in the DDPs offered by these three platforms.
-Note that although Instagram has cookies field it just shows the last four digits of the cookie name and no other details.
-Information Type TikTok Instagram YouTube Minimum expected fields
-User’s usage Watch Y N Y Content Id, ts
-Search Y Y Y Query term, ts
-Comment Y Y Y Comment text, Content Id, ts
-Like Y Y $Y^{g}$ Content Id, ts
-Messages Y Y NA Content, User Id, ts
-Save/Favourite Y Y Y Content Id, ts
-Share(In-app) Y Y NA Content Id, ts
-Share(Across-app) Y N N Content Id, ts
-Interests/Topics Y Y N List of topics
-Time spent $N^{*}$ $N^{*}$ $N^{*}$ Duration/frequency
-User’s content Media Y Y Y Media file/URL
-Text details Y Y Y title
-Location Y Y Y Some place identifiers
-Date time Y Y Y ts
-Device - Y N Device model, OS
-Other user interactions $N^{*}$ $N^{*}$ $N^{*}$ Likes, Comments
-Personal details Account details Y Y $Y^{g}$ Username, DOB, Email, Profile photo
-Connections Y Y Y Username, ts
-Login history Y Y $Y^{g}$ IP, ts
-Current devices Y Y $Y^{g}$ User agent
-Current camera NA Y NA Version/type
-Location Y Y $Y^{g}$ Place identifiers
-Account changes - Y N Type, Old, New values, ts
-Advertisements Ads viewed N N Y Content Id, ts
-Personlization ${N^{*}}$ N N Reasons why the ad was shown
-Access to your data N Y N Which and how (in store visit etc.)
-Miscellaneous Off-platform Y Y N Platform, ts, activity
-Link history - Y - Link, ts
-Cookies - Y - -
-Table 5: Data Transparency: Overview of collected and shared information.
-${N^{*}}$ denotes the details are found in the app but not in the GDPR dump.
-${Y^{g}}$ denotes the details are found in google’s DDP, but not in YouTube’s DDP.
-(ts-TimeStamp)
-Appendix B Survey questions, prompts, and responses from an LLM.
-We included the following in the supplementary material [^6] https://zenodo.org/records/14721885 :
-1. Survey questions used for Section 4.
-2. Overall DDPs used for evaluation in Section 4.
-2. Prompts provided to LLM and responses from LLM used in Section 5.
-4. Survey question used for Section 5.
-Appendix C Comprehensibility of DDPs
-Table 6 presents the distribution of the participants recruited for the survey in Section 4.
-Attribute Type Count Percentage
-Gender Male 228 57.0
-Female 161 40.3
-Other 9 2.3
-Prefer not to say 2 0.5
-Age 18-25 95 23.7
-25-30 104 26.0
-30-40 114 28.5
-40+ 87 21.8
-Country France 100 25.0
-Germany 100 25.0
-Spain 100 25.0
-Italy 100 25.0
-Table 6: Distribution of participants in survey reported in Section 4 based on their self-reported gender, age, and country of residence.
-C.1 Evaluation for the requirements across the platforms
-Table 7 shows the data categories considered for evaluating the comprehensibility of DDPs.
+Figure 5: Figure representing how the watch history from the three platforms was displayed to users for comparison and to evaluate the four properties: conciseness, clear and plain language, intelligibility, and transparency.
 Category What does it mean?
@@
 User’s content List of contents uploaded or posted by the user on the platform.
-Table 7: A list of categories of data that were considered in the adherence evaluation phase of the user survey.
-Figure 8 shows the representation of watch history for the three platforms and sample question format.
-Figure 8: Figure representing how the watch history from the three platforms was displayed to users for comparison and to evaluate the four properties: conciseness, clear and plain language, intelligibility, and transparency.
-Appendix D Improved DDPs comprehensibility
-Table 8 presents the distribution of the participants recruited for the survey in Section 5.
-Attribute Type Count Percentage
-Gender Male 100 50.0
-Female 99 49.5
-Prefer not to say 1 0.5
-Age 18-25 54 27.0
-25-30 57 28.5
-30-40 57 28.5
-40+ 32 16.0
-Country France 50 25.0
-Germany 50 25.0
-Spain 50 25.0
-Italy 50 25.0
-Table 8: Participant distribution in the survey reported in Section 5.2 based on their self-reported gender, age, and country.
+TABLE VI: A list of categories of data that were considered in the adherence evaluation phase of the user survey.
+Survey setup: In the final part of the survey, we presented our participants with the information shared in the DDPs of TikTok, Instagram and YouTube in an anonymized and randomized format as shown in Figure 5.
+We showed different categories of content and the whole DDP to the participants.
+Table VI summarizes the categories that we evaluated.
+[^3] Note that YouTube does not provide like history, autofill, location, and off-platform activity.
+Hence, for these categories, we asked our participants to evaluate between Instagram and TikTok.
+For each category, we showed participants the content shared by the three platforms side by side and asked their preference of data representation they found to be more concise/intelligible/transparent/in clear and plain language.
+We posed this question as a multiple-choice grid (See Figure 5), so each participant had to provide an ordered preference.
+In this context, first preference means the most preferred choice of data representation, whereas the third preference means their least preferred choice.
+Note that the final part of the survey consisted of 14 (13 categories + overall DDP) $\times$4 evaluations where participants gave us an ordered preference.
+Hence, to avoid participant fatigue, we asked half of our participants to evaluate intelligibility and transparency and asked the other half to evaluate conciseness and language properties.
+Hence, 200 participants answered each of the evaluation questions.
+Next, we report our findings on the adherence evaluation.
+Language: Figure 6 shows the percentage of first preference votes elicited by the representation of each of the platforms for the four requirements that we surveyed.
+In the evaluation of clarity and simplicity of language, Instagram won the first preference votes in eleven out of the thirteen data categories (Figure 6(a)).
+In contrast, for the remaining two categories - search and watch history - participants preferred the language in which YouTube presents it.
+Apart from the thirteen data categories, we also asked participants to evaluate the overall DDP from the three platforms.
+55% of the participants preferred the language of Instagram’s DDP.
+TikTok and YouTube DDPs are preferred by 25% and 20% of the participants, respectively.
+Intelligibility: Similar to Language, Instagram won the first preference vote for intelligibility in eleven out of the thirteen categories evaluated (Figure 6(b)).
+In all categories, except for devices and location categories, more than 75% of the participants preferred Instagram’s data representation to be the most intelligible.
+For the evaluation of the overall DDP, 63% of the participants identified the Instagram DDP to be the most intelligible.
+TikTok and YouTube were chosen as the first preference by a considerably smaller proportion of our participants, nearly 18% each.
+Transparency: YouTube won the first preference for transparency requirement in five categories, where over 70% of the participants chose it for watch history, search history, and login history, followed by devices and content with 55% and 46%, respectively For the remaining eight out of the thirteen categories, Instagram won the first preference of the participants in a large percentages.
+For the overall DDP, Instagram is preferred more compared to the other two platforms, with nearly 46% of the participants choosing it to be the most transparent.
+YouTube and Tiktok are preferred by 35% and 19% participants respectively.
+Conciseness: Based on Figure 6(d), we found that in nine out of thirteen categories of data that we surveyed, respondents preferred TikTok’s data representation as the most concise, whereas Instagram’s data representation is found to be the most preferred for the remaining four categories.
+While TikTok’s representation is preferred in the location, autofill, and off-platform categories ($\geq$60% first preference votes), Instagram’s representation is preferred more for likes history and contents ($\geq$65% first preference votes).
+While evaluating the overall DDP, 43% of the participants found Instagram’s DDP to be more concise, followed by Tiktok and YouTube with 35% and 22%, respectively.
+Accessibility: We excluded accessibility from the survey evaluation because it pertains to getting access to request and download DDPs more than to their representations.
+Since it is the most objective requirement for evaluation, we note down the observed differences across platforms below.
+In terms of ease of request, from the content page, one needs to click 6, 6, and 10 times on TikTok, YouTube, and Instagram, respectively, to be able to request the data in the human-readable format.
+Notice that, while on Instagram, a user needs to click more, the platform provides many options to the end-user, such as the time duration.
+Based on our anecdotal observations, the turnaround time for TikTok upon request was instantaneous during this study.
+For Instagram and YouTube, it often takes 10 to 15 minutes for the DDP to become available.
+Thus, all the turnaround times are within the GDPR-prescribed maximum duration of one month.
+One can download one’s data on TikTok and Instagram from both their applications and web versions.
+On the other hand, YouTube shares the DDP by sending an e-mail with a URL to it.
+All these platforms provide DDPs in a compressed format, which makes it harder for users to open them on a smartphone.
+At the same time, a user needs to have a browser or text reader to open the HTML (YouTube and Instagram) and TXT (TikTok) DDPs.
+In summary, while the data request and download are relatively accessible in the current setting, reading the data is still complicated for the average end user.
+At this point, we would also like to note that we are not aware of any special provisions that platforms may have for people with visual impairments to access the data.
+Variations across countries: For the survey based evaluation, we do not observe any qualitative variation across the four countries for three requirements – language, intelligibility, and transparency.
+However, for conciseness, participants from Germany predominantly preferred Instagram, with over 50% of the participants selecting it as their first choice across most categories.
+In contrast, participants from other countries selected TikTok as having the most concise representation for most of the categories.
+Upon a manual inspection of the interpretation of conciseness, we found that, except participants from Germany, most other users interpret conciseness as: (1) minimizing document length, (2) avoiding information overload etc.
+This difference in interpretation is reflected in how our participants evaluated the data representations.
+(a) Clear and Plain Language
+(b) Intelligible
+(c) Transparency
+(d) Conciseness
+Figure 6: Percentage of first preferences across all categories and for the entire DDP.
+Instagram is the top choice for most categories across all requirements, except for conciseness.
+For watch and search histories, YouTube is the most preferred platform in transparency, intelligibility, and language.
+Important takeaways:
+We found participants’ interpretation of the requirements agree with that of European Data Protection Board.
+However, conciseness and transparency requirements of the GDPR are inherently at odds with each other.
+People assess Instagram’s current data representation to be most comprehensible followed by that of YouTube.
+6 Improving comprehensibility of DDPs
+While the previous section uncover a number of shortcomings in comprehensibility of the DDPs, in this section, we present some recommendations to improve their comprehensibility.
+To this end, our proposed recommendations are two pronged (a) improving data representation within DDPs, (b) improving the presentation of DDPs.
+6.1 Improving data representations within DDPs
+Figure 7: Comparison of ‘watch history’ in our proposal and that of YouTube’s current DDP implementation.
+(a) Clear and Plain Language
+(b) Intelligible
+(c) Transparency
+(d) Conciseness
+Figure 8: Percentage of votes for the proposal and the earlier winning platform across different categories for each requirement.
+Our proposal emerged as the top choice in all fields for the language and intelligible requirements.
+To improve the data representation within DDP, we deploy an AI assisted human-in-the-loop approach.
+In this approach, we leverage the different interpretations obtained by survey participants and EDPB guidelines and use a large language model (LLM) to come up with an initial data representation.
+Then, we modify these initial representations to propose the final data representation.
+6.1.1 Methodology
+To come up with a recommendation for data representation, we take the following steps.
+Step 1: Summarize the interpretations: For each of the GDPR requirements, we collect the interpretations from our participants and EDPB guidelines [12].
+Digesting such a diverse set of requirements and their expectations, and ironing out data representations that adhere to such requirements and live up to the expectations of different stakeholders is cognitively challenging even for experts.
+Hence, we feed this entire set of interpretations to an LLM – Gemini 2.0 Flash [8]– using the Google AI Studio platform.
+We ask the LLM to generate a set of representative sentences that encapsulate the most prevalent interpretations by summarizing the above responses.
+For reproducibility purposes, we make the prompts we used and the responses available in the GitHub repository.
+Note that in this approach, we include the interpretations of both data subjects and data protection authorities (through the EDPB interpretations).
+Step 2: Generate recommended data field for each data category: In the second step, we pass the representative sentences obtained in Step 1 for each of the requirements to the LLM to obtain the specific data fields for each data category that we should provide to the end user.
+Subsequently, the LLM provides us with a set of recommendations for data fields of information in the queried category in the form of a JSON object (available in the GitHub repository).
+Step 3: Manual inspection by researchers: In the final stage, three researchers (coauthors of this paper) go over these recommendations to assess their quality.
+The primary goal of their assessment is to remove any instance of hallucination in the LLM responses; secondly, they wanted to ensure not to recommend some data fields that will enable platforms to collect any additional personal data.
+Next, given their expertise in understanding data donations in other platforms like Netflix, Prime video etc., we ask them to suggest any additions that were deemed relevant or removal of certain fields when they contained excessive detail in the recommended fields.
+Note that while Step 1 introduces the end users’ and data protection authorities’ interpretations, this step also instills the interpretations of tech experts on other platforms and researchers who want to study the dynamics of platforms using these DDPs.
+Overall these interventions affected changes in 20 out of 85 data fields across all the data categories.
+By taking these steps, we ultimately come up with the final recommendation of improved representation for different data categories.
+6.1.2 Evaluation of the proposed representation
+To evaluate our proposed representation, we conducted another survey among 200 participants (who have not participated in the survey in Section 5) from Germany, France, Spain, and Italy.
+We follow the participant recruitment strategy as elaborated in Section 5.
+Table VIII in Appendix B shows the user demographics.
+During the survey, we showed the participants two representations for the evaluated categories: (a) proposed representation: which we generate from the steps mentioned in Section 6.1.1, (b) winner representation: for each data category, the representation among Instagram/TikTok/YouTube which won the first preference along more requirements (example, as shown in Figure 7).
+We randomly presented each pair of anonymized representations to the participants and asked them to select the representations that they find to satisfy the requirements more.
+Observations: We compare the representations of the data category fields of our proposal and that of the best among the existing short-format video platforms for 13 different categories across four GDPR requirements.
+[^4] We are not considering personal information, as there were no suggestions for adding or removing any fields.
+Instead, we compare activity summary, which is provided only by TikTok.
+Figure 8 shows the percentage of votes for the two representations obtained in the 52 evaluations (13 data categories $\times$ 4 requirements), that we surveyed.
+We observe that the participants preferred our proposed representation in as many as 44 categories.
+However, their preferences vary for different requirements.
+Language and intelligibility: For clarity and simplicity of language and intelligibility of the representation, our proposal gets upwards of 67% of the votes in most of the categories , suggesting our proposed representation is both intelligible and in clear and plain language.
+Transparency: Our proposal is voted to be more transparent than the best of the existing representations in 11 out of the 13 categories.
+Apart from search history, in all the other winning 10 categories, our representation gets upwards of 75% of the votes.
+However, in the autofill information category, Instagram’s data representation gets nearly 93% of the votes from our survey participants.
+While Instagram’s representation contain many sensitive information, our recommendation contained a subset of them needing less privacy-intrusive data gathering.
+Conciseness: Notice that in Section 5.2, we mentioned conciseness and transparency are mutually competing requirements.
+We observe the same conflict in the results of this survey.
+Our proposal was preferred to be the most concise representation in only 7 out of the 13 categories.
+In watch history, search history, login history, and device details – our proposed representation was found to be more concise by upwards of 75% of the participants.
+Also, similarly to the survey presented in Section 5.3, participants from Germany have different preferences compared to those from other countries for conciseness.
+Except for autofill information, over 66% of German participants preferred our proposed representation in all other categories.
+The primary objective of this section has been to convey how one can come up with more effective data representations by considering the interpretations of different stakeholders.
+Our observation potentially indicates that currently platforms are probably operating in silos without much deliberation and sincere effort into their implementation.
+At the same time, because of the mutually conflicting nature of the requirements, it is impossible to satisfy all the requirements with any single data representation.
+6.2 Improving the presentation of DDPs
+While the recommended representation improves the content within DDP, due to the mutually conflicting nature of the requirements (e.g., conciseness vs.
+transparency), it is impossible to satisfy all the requirements with any single data representation.
+Therefore, a second layer of improved presentation, in the form of a user-centric dashboard, is needed to reconcile such trade-offs.
+Along with improving along other requirements, this approach provides users the autonomy to choose their desired level of transparency.
+6.2.1 Methodology
+We take the following steps to improve the comprehensibility of DDPs to end users.
+Figure 9: ① shows how YouTube currently shares watch history.
+② illustrates how our approach addresses some of the transparency expectations of our survey participants.
+③, ④, and ⑤ depict our proposed concise, raw data, and transparent views respectively.
+Refer to Figure 11, 12 in Appendix C for TikTok and Instagram respectively.
+Note that this is partial view.
+The full view of YouTube Dashboard can be accessed in the GitHub repository.
+Step 1: Understanding user expectations from DDPs.
+To understand user expectations from DDPs, we conducted a structured survey focused on the data categories analyzed in Section 5.
+A total of 120 participants were recruited from four EU countries – Germany, France, Italy, and Spain – with 40 participants assigned to each platform.
+Demographic details of the participants are presented in Table VIII in Appendix B.
+In total, there were 1,569 responses regarding expectations across all categories and the three platforms.
+Using BERTopic, we clustered these responses, which revealed a broad spectrum of interests.
+For instance, for YouTube browsing history (see Figure 9 ①), questions ranged from high-level insights – such as identifying the most frequently viewed channels or determining periods of peak activity – to low-level clarifications, including interpretations of specific data fields (e.g., understanding what “product” refers to in a given context).
+Further, several recurring concerns emerged across participants, including questions about the purpose of data collection, data access, retention duration, and control over data collection.
+Low-level clarifications accounted for 8.85% of the responses, while transparency-related questions (purpose, access, retention, control) contributed 17.5%.
+The remaining responses were high-level insights which are summarized in Table VII.
+Category Sample expectations Percentage
+Watch history Which videos I watched the most? How long did I watch the video? 17.8
+Search history How many searches do I do per day or per session? What did I search? 11.7
+Comments How often do I comment? On what post did I comment? 7.3
+Saved How many saved posts do I have? Do I revisit my saved posts? 3.6
+Connections I would like to know when someone started following me. I’d like to know when I subscribed to this channel. 8.6
+Devices What devices did I use? From what device and the % of use per device? 4.8
+Login history What times did I most log in? How long did the session last for that login? 6.3
+Like history How many likes I did? Whose content do I like more? 5.0
+TABLE VII: User expectations from different data categories in the DDPs.
+Step 2: Designing a user-centric dashboard.
+Based on the insights gathered from the survey, we developed an interactive and user-centric dashboard that empowers individuals to explore their data in a manner aligned with their preferences.
+The dashboard offers three key modes of interaction: (1) a concise view: which presents high-level insights in a simplified and digestible format, (2) a raw data view: which presents the data exactly as provided by the platform, but with clearer and simpler language for better understanding; and (3) a transparency view: which includes field-level explanations that clarify the intelligibility and utility of each attribute.
+Also, we answer the concerns such as the purpose of data collection, data retention duration, and who has access to the data by collecting relevant information from platforms’ privacy pages.
+Figure 9 shows the different types of views we proposed for YouTube watch history.
+Step 3: Delivering a browser extension tool.
+We package it in the form of a browser extension: Know Your Data.
+The code is made publicly available.
+In addition, the extension also automates the process of requesting and downloading data from platforms thereby improving accessibility.
+The browser extension was implemented using React.js [26] for the frontend interface, Vite [34] as the build tool, and Recharts [27] to enable interactive visualizations.
+Performance profiling using Chrome’s task manager indicated CPU utilization below 2% when idle, averaging approximately 25% during typical processing, with transient peaks up to 110% during word cloud generation.
+This browser extension-based solution has key advantages in the following aspects: (1) Privacy: Significantly more privacy-preserving as sensitive data never leaves the user’s control or device, (2) Security: Reduces the risk of data breaches on our end, as we never handle or store the user’s DDP on our servers, (3) Trust: Users are generally more comfortable with extensions that process data locally, (4) GDPR compliance: Much simpler from a GDPR perspective because we are not acting as a ‘data processor’ in the sense of receiving and storing data on our infrastructure.
+Generalizability: The methods we proposed for improving the comprehensibility of the DDPs are generalizable to other platforms.
+To improve the representation, we first summarized the user interpretations which are platform independent and generalize as-is across platforms.
+We then identified platform specific data categories and incorporated them into the LLM prompt.
+These categories can be easily extracted by inspecting the DDP of each platform, but they generalize to DDPs of all users of a given platform.
+Finally, researchers who have expertize in understanding DPPs manually reviewed the LLM-generated responses.
+The corrections required were minimal, as the LLM was already able to produce sufficiently accurate DDP representations.
+For improving the presentation of DDPs, we surveyed users to know their expectations from the DDPs.
+These expectations can be generalized to other platforms, particularly transparency-related questions, whereas the high-level insights and clarifications depend on the type and granularity of data each platform collects.
+Important takeaways:
+Our recommended data representation, that considers interpretation of different stakeholders, was favored by surveyed participants in 44 out of the 52 evaluations conducted across various pairs of data categories and requirements.
+Simply improving data representation does not improve adherence to Article 12 requirements of the GDPR since some of the requirements are at odds with each other.
+Understanding user expectations and giving them the autonomy may be an effective way to reconcile the trade-offs between the requirements.
+Our proposed Know Your Data extension is one such operationalization.
+7 Ethical considerations
+We acquired approval from the Ethical Review Board (ERB) of our university to conduct this research.
+We obtained explicit consent of each participants before they participate in our surveys.
+Importantly, we refrained from collecting any PII from our participants during the surveys or from their DDPs.
+Although DDPs may contain PII, they are usually stored within specific keys or files.
+To refrain from collecting them, we inspected the DDPs of our own accounts on the three platforms and collected a predefined list of keys and files that contain any PII.
+Next, we developed scripts that remove these specific keys and files inside a DDP on the user’s end, before collecting the modified DDP to our server.
+Furthermore, based on our ERB, we would not share the donated DDPs with any third parties, and we will delete them within 3 years of completion of our study.
+Finally, in Section 4, we used bot accounts that interacted with publicly available content on the platforms.
+We further took steps to minimize the consequences of running these bot accounts.
+Specifically, we distributed the overhead of our experiment over multiple short browsing sessions.
+Furthermore, even though the Terms of Service of the studied platforms prohibit the usage of automated bots, we believe that the benefits of our study outweigh any potential associated unintended consequences.
+8 Concluding discussion
+Summary: To the best of our knowledge, this is the first systematic audit of the implementation of GDPR’s Right of Access from the lens of reliability and comprehensibility.
+By auditing the implementations across platforms offering similar services, we uncover a number of platform-specific and systemic instances of non-compliance.
+While periodic audits can ensure reliability compliance, to improve the comprehensibility compliance we also recommend a two-layered approach: (a) an AI assisted human-in-the-loop approach to improve DDP representation, (b) a Know Your Data extension to improve DDP presentation.
+Our recommendations not only aim at striking a better balance between GDPR’s conflicting comprehensibility requirements, but also provide more autonomy to users to choose their desired level of transparency in data visualization.
+Limitations of the current work: Our study, while comprehensive, has several limitations.
+Our survey participants, recruited via Prolific, may not fully represent the broader spectrum of digital literacy or demographic diversity of the European Union.
+However, note that our work is comparative by nature.
+Hence, we do not expect the choice of participants will impact the findings presented in the paper.
+Further, our study focuses on data shared by platforms upon request and not the data that is collected.
+Consequently, our findings on transparency of data provided are based solely on what platforms choose to share in DDPs, which might not reflect the entirety of data they process.
+Including this information may further enrich the quality of the work.
+Recommendations for stakeholders: First, platforms should improve the completeness, correctness and consistency in data granularity, retention periods, and included metadata across related data categories for a more effective and user-centric implementation.
+Furthermore, platforms should include the purpose of processing, retention periods, and third-party recipients of collected data.
+Such inclusion can not only improve the legal compliance but also match the user expectations from the DDPs (as observed in our survey for designing the dashboard).
+Finally, as shown in Section 6, rather than deciding on behalf of users, platforms should design solutions that empower users with more autonomy to make informed decisions.
+While our findings of non-compliance provide regulators with starting points to investigate the platforms, our observation on conflicting requirements should nudge regulators to think through the requirements they mandate for implementations.
+In fact, such rights will only be useful if data protection authorities provide more detailed technical specifications and/or standardization for the implementation and enforcement of these seemingly conflicting requirements.
+Finally, this study demonstrates the value of employing computer science research and scalable audit strategies to effectively monitor and enforce GDPR compliance.
+Acknowledgments
+We extend our sincere gratitude to Prof.
+Christoph Engel (Max Planck Institute for research on Collective Goods) and Prof.
+Cristoph Sorge (University of Saarland) for their valuable feedback.
+Additionally, Ingmar Weber is supported by funding from the Alexander von Humboldt Foundation and its founder, the Federal Ministry of Education and Research (Bundesministerium für Bildung und Forschung).
+Appendix A Reliability of DDPs
+Figure 10 represents the plots that show the CDF of data duration provided to users for the like history for the two platforms - Instagram and TikTok.
+We do not observe any clusters, suggesting that these platforms provide the like history more consistently to the end users.
+Appendix B User demographics for different surveys
+Table VIII presents the distribution of the participants recruited in Section 4, Section 5, Section 6.1 and Section 6.2.
+Appendix C Improving DDPs comprehensibility
+Figure 11 and Figure 12 are the partial screenshots of the dashboard for browsing history on TikTok and Instagram as proposed in our extension Know Your Data.